### v3

# Intuitive Surgical SurgToolLoc Challenge Results: 2022-2023

Abstract
Robotic assisted (RA) surgery promises to transform surgical intervention. Intuitive Surgical is committed to fostering these changes and the machine learning models and algorithms that will enable them. With these goals in mind we have invited the surgical data science community to participate in a yearly competition hosted through the Medical Imaging Computing and Computer Assisted Interventions (MICCAI) conference. With varying changes from year to year, we have challenged the community to solve difficult machine learning problems in the context of advanced RA applications. Here we document the results of these challenges, focusing on surgical tool localization (SurgToolLoc). The publicly released dataset that accompanies these challenges is detailed in a separate paper https://arxiv.org/abs/2501.09209 [1].
1 Introduction
Robotic-assisted (RA) surgeries become more prevalent with each passing year [2, 3]. Importantly, RA surgeries also generate rich streams of data, e.g. video recordings, tool kinematics, and events gathered from the robotic system, that lend themselves to quantitative analyses. Pursuits such as the quantification of surgical performance, autonomous analysis of procedures, surgical skill evaluation, AI-guided surgical planning, and surgical data science, in general, can all exploit this rich source of data [4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. However, despite the ability to collect large amounts of it, annotating it to provide meaningful context remains a major bottleneck in the field.
Acquiring annotations for surgical data is particularly challenging as it requires personnel with in-depth domain knowledge. What’s more, relying solely on human annotators is costly, time-consuming, and subject to inter-rater variability. Yet RA procedures generate numerous, potentially instructive, labels that can be harvested automatically from the system. For example, each instrument install (or uninstall) elicits events that record the tool type and the install (uninstall) time. Similarly, commands to these tools, such as the use of energy, elicit related events. While events such as this are informative, they are not always representative of the tools visible in the surgical field, let alone their spatial locations. Machine learning techniques that could reliably exploit these events, by interpreting them as noisy, weak labels for training, would prove extremely valuable. While algorithms for this class of problem exist, there is a real need for advancements in their accuracy and robustness.
Machine learning (ML) challenges –also known as ML competitions– have become an integral part of many top computer vision and machine learning conferences [14]. These competitions aim to address specific problems within various fields where the organizers provide the required data set and/or labels. In the surgical domain specifically, EndoVis [^1] https://opencas.dkfz.de/endovis/ is an annual challenge held at the international Conference on Medical Image Computing and Computer Assisted Intervention MICCAI [^2] https://miccai.org/. This challenge contains multiple sub-challenges targeting problems like tissue and instrument tracking and segmentation, 3D scene reconstruction, action recognition, and surgical skills assessment [15, 16, 17, 18, 19]. Following suit, our team has organized a yearly challenge since 2020.
With varying differences, all our challenges have focused on problems related to RA surgical intelligence. Beginning in the year 2022, however, we have focused on the same problem of weakly supervised learning of surgical tool localization (SurgToolLoc). For these challenges we provided participating teams with an extensive set of endoscopic video data. Alongside these videos, the teams received noisy annotations for the surgical tools present in each video frame, in the form of tool presence information. The ultimate goal of these challenges was to utilize automatically extracted tool labels to build models that can classify and localize surgical instrumentation in a hidden test set. In addition, each year we have posed a unique secondary challenge, that also utilizes the same video data. However, the main objective of our repeated challenge is to facilitate year over year advances in the ML solutions to the weakly supervised problem of tool classification and localization.
Here we document the results of each year’s challenge. We begin with an overview of the data set made available to the teams. Then we present each year’s challenge as a separate chapter. We end with some general observations, outlining what strategies have proved most successful, and what we believe are the challenges moving forward. Finally, we note that in the interest of encouraging the surgical data science community to continue working on similar problems, the training data used in these challenges has be released publicly in a separate publication. To the best of our knowledge, this is the largest publicly available surgical data set to date. It is our firm hope that this ongoing challenge and the accompanying data sets become touchstones for the broader field of surgical data science.
2 Data Overview
For this continuing challenge a large dataset of surgical videos and accompanying annotations and labels was generated. Video and system data were harvested from surgical training exercises performed on the da Vinci robotic system. During these exercises, surgical trainees (and some experts) performed a series of standardized activities on porcine tissue, such as dissecting, suturing, and cauterizing. The captured video displays the trainee’s view from the surgeon’s console, although it is only one of the two video streams from the stereoscopic imaging system.
The training exercises that generated the data were performed on two distinct animal modalities, either an anesthetized living pig, or using organs harvested from a pig. This distinction leads to differences in the general appearance of the resulting video frames, and those of the former modality are generally redder in tone. The anatomy present in the videos was not relevant to this challenge.
To obtain training labels system data was harvested while the exercises were performed, and a team of trained clinical experts annotated the videos. The system data provided the tools installed in the robot. The clinical experts annotated the videos, segmenting them into surgical steps. The challenges hosted in 2022 and 2023 drew from this dataset. A separate paper that accompanies the dataset provides more details along with download link [1].
3 Results and methods from the MICCAI 2022 SurgToolLoc challenge
3.1 SurgToolLoc 2022 Challenge Description
3.1.1 Overview
The sub-challenge was arranged as part of the Endoscopic Vision Challenge at MICCAI 2022. The challenge had two separate categories open to the participating teams (Figure 1). Category 1 focused on identifying the presence of surgical tools in video frames, while Category 2 challenge was designed in order to identify both those tools present and their locations in the corresponding video frames.
Both categories required the development of weakly supervised algorithms. The first category relied on supervised learning for object detection with weak/noisy labels. The second category required weakly supervised object detection and localization with the same labels. Teams were allowed to participate in either challenge or both.
3.1.2 Category 1: Surgical Tool Classification
This category required the teams to submit a model that could classify all the tools present (up to 3 surgical instruments) within 30-second video clips. This required the teams to train a multi-label model to classify the 14 possible tools. A data set of 24,695 video clips was provided for training. The accompanying labels indicated which robotic tools were present in each clip. For technical reasons, a portion of the labels was missing or incorrectly assigned.
Submissions were evaluated on a hidden test data set of surgical videos through the Grand Challenge’s automated docker submission and evaluation system. Teams were ranked by their model’s average F1 score across all classes.
3.1.3 Category 2: Surgical Tool Classification & Detection
This category required the models to perform the same classification problem as before while also localizing the identified tools with bounding boxes for each frame. As in Category 1, the models localized the tools present within a set of 30-second video clips.
Figure 1: Overview of challenge categories 1 and 2
3.1.4 Training Data
Video and system data were harvested from surgical training exercises performed on the da Vinci robotic system (see Section 2). From this dataset 30-second video clips were sampled. Across all the video clips, 14 tools were installed and visible. They were: needle driver, cadiere forceps, prograsp forceps, monopolar curved scissors, bipolar forceps, stapler, force bipolar, vessel sealer, permanent cautery hook/spatula, clip applier, tip-up fenestrated grasper, suction irrigator, bipolar dissector, and grasping retractor.
From the data collected, 24,695 video clips from a set of 11 unique training exercises were made available for training purposes. Each clip was 30 seconds long and captured at 60 FPS with a resolution of 720p (1280 x 720) from one channel of the binocular endoscope (Figure 2). For the extent of each clip, there were up to three robotic surgical tools simultaneously installed within the surgical field. For each video clip within the training set, we provided the corresponding tool presence labels that indicated which robotic tools are installed. Since the tools are installed for the entire duration of each video clip, only one label per video clip was provided. Notably, for some clips, or portions of clips, tools were obscured or otherwise temporarily not visible. Moreover, the instrument distribution among the data was highly imbalanced. As such, the tool presence labels were noisy indicators of the presence of a surgical instrument in the video frames.
Presence labels were a list of four strings corresponding to the tools present in the robot’s four arms [USM1, USM2, USM3, USM4], e.g, [’cadiere forceps’, ’nan’, ’force bipolar’, ’grasping retractor], where the arm with the camera installed is labeled as ”nan.” The four arms are usually (though not necessarily) installed from left to right (i.e. USM1 is the leftmost arm, and USM4 is the rightmost arm). Figure 2 shows an example of the .csv file with the presence labels. This training data set has also been released publicly and can be accessed at https://console.cloud.google.com/storage/browser/isi-surgtoolloc-2022.
Finally, we note that the teams were also allowed to pre-train their models on publicly available surgical datasets such as Cholec80 [20] and M2CAI [21], and comprehensive datasets such as ImageNet [22].
Figure 2: Sample frames with presence labels (left) and a snapshot of the labels CSV file (right)
3.1.5 Testing Data
The testing dataset consisted of 93 video clips taken from surgical training exercises—similar to the training set—using the da Vinci robotic system. The length of each video clip was variable-mean(SD), in seconds: 747.31(579.92)-, but had the same resolution and was sub-sampled to 1 frame-per-second (FPS) for inference purposes. The testing dataset had the tool presence labels in the same way as for the training set but also combined with annotated bounding boxes around the instruments. These annotations were generated by a crowd (30) of experienced annotators. Two rounds of annotation reviews were conducted to ensure quality. It is important to note that the clevis of each robotic surgical tool was considered the ground truth for most tools as Figure 3 shows. However, there are following exceptions to this rule: If a tool’s clevis is not well defined or if the tool is especially large and the clevis does not always show in the field of view, then the bounding box includes the surgical tip as well - e.g monocular curve scissor as shown in the purple-colored bounding box on the left of Figure 3. Moreover, using the information available in the UI to make predictions was not allowed. To enforce this, the area comprising the UI interface in the test set was blurred, as seen in Figure 3.
Figure 3: Sample frames with testing labels. The UI interface was blurred to avoid its embedded information (description of the tools) being used for predictions.
3.1.6 Submission Process
The challenge was hosted as a “Type 2 (T2) challenge” on the grand-challenge (GC) platform333grand-challenge.org. T2 challenges on GC allow for the automated submission of containers, while the evaluation process for these submissions is executed via a standardized evaluation protocol to ensure the reproducibility of the algorithm. This modality also allows for each algorithm to be evaluated on a private test dataset that is hidden from the teams.
Each team developed its own algorithm following the guidelines provided by the GC platform [^4] https://grand-challenge.org/documentation/algorithms/ and the organizing team of the challenge. Every submitted algorithm was configured to receive as input either a preliminary test dataset containing two videos for debugging purposes or the final test dataset described in 4.1.3. The algorithm’s output had to follow a standardized file format for each category. For category 1 -surgical tool classification-, the output file had to generate a list of dictionaries containing information on the tools present in each frame of the input video in a boolean format. For category 2 - Surgical tool classification and localization-, the algorithm had to output a dictionary containing the set of tools detected in each frame with its corresponding bounding box cartesian coordinates (x, y). Details on the creation of the algorithms, samples of input data, and the required format for output files were provided to the participating teams  [^5] https://surgtoolloc.grand-challenge.org/submission/.
Following each evaluation, the platform’s embedded leaderboard displayed each team’s submission scores. Then, the teams were asked to submit a final report with their methodology and results when evaluating the algorithm with the hidden test dataset. Each report is reproduced in this paper in section 3.2.
3.2 Team Submissions
For this challenge, there were a total of 79 teams that had shown interest in participating and downloading the dataset. In the end, there were 17 teams that made full submissions in Category 1 whereas 11 made full submissions for Category 2. All teams that participated in Category 2 also participated in Category 1. A table showing the participating teams is presented below (Table 1), while Table 2 summarizes the methodologies employed by each team. What follows in its section are the methodological details of the participating teams. Please note that these sub-sections were written by the participating teams so a difference in writing style and content will be observed.
Table 1: Team affiliations and challenge categories.
Team # Team name Institution Country Category Report
1 HRI_MV Hikvision Research Institute China 1 & 2 Y
2 HKMV South China University of China China 1 & 2 Y
3 NVIDIA NVIDIA USA 1 & 2 Y
4 ANL-Surg Argonne National Lab USA 1 & 2 Y
5 HVRL Keio University Japan 1 & 2 Y
6 SK Muroran Institute of Technology, Niigata University of Health and Welfare, Konica Minolta Japan 1 & 2 Y
7 TeamZERO University of West of England Bristol, University of Glasgow, University of Bristol UK 1 & 2 N
8 VANDY-VISE Vanderbilt University USA 1 & 2 Y
9 UKE University Medical Center Hamburg Germany 1 only Y
10 Gatech Georgia Institute of Technology USA 1 & 2 N
11 ITeM Furtwangen University, University of Lipzig Germany 1 & 2 Y
12 MM Chinese University of Hong Kong, National University of Singapore, University College London Hong Kong, Singapore, UK 1 only Y
13 BioMedIA Muhmammad bin Zayed University of Artificial Intelligence UAE 1 only Y
14 WhiteBox The University of Tokyo Japan 1 only Y
15 CAMMA University of Strasbourg France 1 only Y
16 Vision_HK Tianjin University China 1 & 2 N
17 lsgroup Xiamen University China 1 only N
Table 2: Summary of methodologies
Team Name Architecture Backbone Data Preprocessing Pretrain Image Augmentation Use Additional Data Loss Output HKMV Faster-RCNN efficient net-b3 Endovis17, Endovis18 N/A Bounding box generated from segmentation mask tool bounding box SK Multiplicity Feature fusion ResNet-50 Resize Segmentation data Yes N/A tool class, tool bounding box NVIDIA Hybrid (ConvNext-tiny, EfficientNet-B4, Ensemble), YoloV5 Resize, crop Yes Pseudo label and manual bounding box tool class, tool bounding box WhiteBox CNN ResNet-18 Resize, crop Asymmetric Focal Loss tool class HVRL Hybrid (Swam Transformer, EfficientNet-V2, ResNeXt) Resize, crop ImageNet weights, Instagram weights Yes binary cross-entropy tool class, heatmap, tool bounding box MM Hybrid (ResNet50, ViT-Batch16) Resize Yes Asymmetric Focal Loss tool class CAMMA Spatial Attention network (SANet) ResNet-18 Resize ImageNet weights Yes weighted sigmoid cross entropy tool class, attention map UKE DINO-based ViT ViT Resize ImageNet weights Yes tool class, attention map, tool bounding box BioMedIA YOLOv5, ConvNext ResNet 50 Oversampling, UI removal Cholec80, EndoVis’15, CholecSeg8k Yes Cholec80, EndoVis’15, CholecSeg8k datasets focal loss, box regression loss, objectness loss tool class, tool bounding box ANL-Surg Ensemble (YOLOv5, ConvNext); Detectron2 Resize, crop MaskRCNN weights - Segmentation labels from MICCAI17 and MICCAI18 tool class, tool bounding box VANDY-VISE Query2Label ResNet, ViT - Asymmetric loss tool class, tool bounding box ITeM CNN + Squeeze and Excitation (SE) module ResNet50 Cholec80 - binary cross-entropy, softmax cross-entropy tool class, tool bounding box
3.3 Hikvision Research Institute - Team HRI_MV
Team Member: Quan He
The HRI_MV team made full use of the characteristics of the video data and adopt an object-tracking scheme with a semantic segmentation network to obtain the pseudo tags of the frames. In particular, an IOU scoring with the dual model cross-validation mechanism was employed to choose more reliable pseudo tags. Using the obtained pseudo tags, a detector was trained and integrated into a pseudo tag using a Kalman filter, iteratively updating the models.
In traditional instrument target detection, the whole object is considered as the ground truth (GT) and included in the detection frame, while in surgical instrument detection, only the clevis is considered as GT. The difficulty is that the appearance of the clevis of different devices is relatively similar. When human eyes distinguish different devices, they are actually mainly distinguished according to the clamps of the devices. To solve this problem, the ROI proposal box output by the RPN network was modified: for the regression head, the proposal box output by RPN was used, while for the classification head, the proposal box was expanded to obtain more information for device classification.
Finally, the integration algorithm was adapted to use the detection frame output from the previous frame for target tracking while combining the results of target detection and target tracking by weighting, which effectively solved the problem of target missing detection.
3.3.1 Method Description
Object Tracking with Cross Supervision:
The HRI_MV team choose to label the frame sequence of the video by using the target tracking method: The first frame of the video was manually labeled, while automatically labeling the subsequent frames using the target tracking algorithm. However, the tracking process sometimes failed. The team found that this was caused by clevis’ small and inconspicuous features, reflection, and other factors. Therefore, a cross-validation method based on segmentation results was proposed.
Siamfc++[23] was chosen for single target tracking (SOT). The advantage of this approach is that the trained SOT model can be used without retraining. Only the first frame of the sequence needed to be labeled so the SOT was able to track the subsequent picture frames effectively. Then, an unet-like segmentation network was trained with the open source data dVRK segmentation dataset, Event-based-2D-tracking-of-articulated-instruments-for-robotic-surgery dataset, EndoVis17 dataset, and EndoVis18 dataset.
The detector was trained with part of pseudo labels and then used to generate new pseudo labels with the tracking output. For frames with an IoU less than the threshold value of 0.2, the generated pseudo label was considered untrustworthy. In this way, a dataset containing about 25000 frames with pseudo labels was obtained.
ROI Expansion:
Cascade RCNN [24] was chosen as a target detection model. In the initial test results, although many instruments were successfully detected, they were misclassified. This is because, in traditional instrument target detection, the whole of the object is considered as GT and included in the detection frame, while in surgical instrument detection, only clevis is considered as GT. Although, the appearance of clevis components of different instruments is relatively similar, and the difference is mainly reflected in the clamp part.
To solve this problem, the ROI proposal box output by the RPN network was modified. For the regression head, the proposal box output by RPN was used, while for the classification head, the proposal box was expanded twice in width and height, respectively, to obtain more information for device classification, as shown in Figure 4.
Figure 4: HRI_MV: The information required for detection and classification is different. Thus, the ROI proposal box output by RPN network was modified.
Figure 5 shows the feature map extracted from the backbone of the proposed model. In the original model, the instrument cadiere forceps was mistakenly considered prograsp Forceps and tip-up fenestrated Grasper, the thermodynamic values in its characteristic diagram are all concentrated in clevis. With the proposed expanded ROI, the cadiere forceps is correctly classified: in its feature map, there is not only a large thermal value distribution at the clevis but also a large thermal value at the clamp, which helps the model to better classify instruments.
Figure 5: HRI_MV: ROI expand makes model identification device
Model ensemble:
The detection frame output from the previous frame algorithm was used for target tracking and integrated into the results of target detection and target tracking via weighting. In this way, as long as the device appears in a certain frame, it can be identified by target tracking in its subsequent frames, solving the problem of target missing detection. The overall framework of our algorithm is shown in Figure 6.
Figure 6: HRI_MV: model ensemble
3.3.2 Preliminary Performance
A test set was established to verify the performance of the detector to evaluate whether more data is needed. The test set was determined by manual review to include each sample category. Roundn means nth times object tracking with the cross supervision method was used to get more data (see Table 3).
Table 3: HRI_MV: Object tracking with the Cross Supervision method was used to get more data.
Method Round1 Round2 Round3 Round4
mean mAP 0.164 0.347 0.489 0.491
By the fourth round, performance has improved slightly. Then, these data were used for the next experiment. The entire dataset was then divided into a training set and test set (7:3) to verify what strategies could be taken to enhance the performance of the model (table 4).
Table 4: HRI_MV: Model Results
Method mean mAP
Base +0.489
+ swin pre-trained +0.052
+ roi expand 2 +0.063
+ ChannelShuffle-RGNShift +0.021
+Fog-Blur-Noise +0.048
+ focallosss +0.002
+cbam +0.033
+DetecoRS pre-trained +0.059
swin pre-trained + roi expand 2 + Fog-Blur-Noise + ChannelShuffle-RGBShift +0.109
detectors pre trained + roi expand 2 + Fog-Blur-Noise + ChannelShuffle-RGBShift +0.111
The mentioned strategies can improve the performance of the model on the current test set; by adopting the ROI expand method, the overall performance was improved by 0.063. Then, the last two models in Table 4 were selected as the models for the ensemble. On the Surgical tool localization leaderboard, the effectiveness of the ensemble scheme was confirmed: the single model reached a score of 0.3985 on the final test set, while the ensemble model reached a score of 0.4077 and ranked first. Since the detector can classify the target, it was also joined with the Surgical tool detection Tasks. However, on the Surgical tool detection leaderboard, The performance of the ensemble model deteriorated when compared to that of the single model, with the former score being 0.7485 and the latter being 0.7341.
3.4 South China University of China - Team HKMV
Team Member: Weiliang Chen
3.4.1 Method Description
In order to obtain an object detection model that can effectively locate surgical tools, the key is how to build a well-labeled training dataset at a low cost. In this method, two public datasets-endovis17 [25] and endovis18 [19]- were used as the primary training dataset. The clevis part in the mask label of the dataset was converted into a bounding box and then used to train an object detection model. This model was then used to infer the images of the competition dataset and then adjust the predicted bounding box slightly. After repeating several operations, a large-scale training dataset can be obtained.
In terms of the model, a two-stage object detection model was built based on Faster R-CNN. EfficientNet-b3 was chosen here, which takes into account both performance and speed, as the backbone to extract image features and add FPN structure to further optimize the feature map.
Dataset:
Two public datasets endovis17 and endovis18 were used to assist us to obtain a primary training dataset. These two data sets provide mask annotations, as shown in Figure 7.
Figure 7: HKMV: Primary training dataset
The clevis part in the mask label of the dataset was converted into a bounding box label and used as the primary training dataset. Since endovis17 and endovis18 do not include all 14 surgical tools, part of the data was added to the training dataset to ensure that the dataset contains all 14 surgical tools. The final set contained 5212 images.
Surgical Tool Localization Algorithm:
  • A Faster R-CNN object detection model was trained on the primary training dataset.
  • EfficientNet-b3, which takes into account both performance and speed, was chosen as the backbone to extract image features and add FPN structure to further optimize the feature map.
  • According to the size of the clevis of surgical tools in the training dataset, four layers of feature maps were included in the model, which were downsampled 8, 16, 32, and 64 times respectively.
Dataset Expansion and Adjustment:
Figure 8: HKMV: Model architecture
In order to increase the size of the dataset, the trained model was used to infer the images of the competition dataset, adding these to the training dataset. A score threshold of 0.7 was employed to filter out the poor bounding box. After several operations, the dataset was expanded to 11035 images. The dataset expansion flow diagram of the algorithm is described in Figure 9.
Figure 9: HKMV: Data expansion scheme
3.4.2 Preliminary Performance
For the preliminary experiments, 25% of the dataset was used as the validation dataset, and only well-labeled images were used in order to accurately evaluate our algorithm. The model structure of EfficientNet-b3 was chosen as the backbone, downsampling 8, 16, 32, and 64 times the feature map. During preliminary experiments on this validation dataset, the model was able to reach a mean mAP of 0.452.
3.5 NVIDIA - Team NVIDIA
Team Members: Bo Liu, David Austin, Yiheng Wang, Michal Futrega, Jean-Francois Puget
This team consists of five NVIDIA employees. Three of them (Jean-Francois, David, and Bo) are members of the Kaggle Grandmasters team, with extensive experience in computer vision machine learning competitions. One member (Yiheng) works on MONAI, an open-source framework for deep learning in healthcare imaging. Another (Michal) works on building GPU-optimized models for NVIDIA’s Deep Learning Examples repository [^6] https://github.com/NVIDIA/DeepLearningExamples. This team’s motivation was to win prizes in both categories using the NVIDIA software stack.
3.5.1 Method Description
Category 1 is a weakly supervised category in the sense that classification labels are given on video level instead of frame level, but test predictions need to be performed on frame level. Since many video-level tools don’t appear in all the frames of the video, the major challenge of Category 1 is to identify frames that contain clean labels. They came up with a unique way to tackle this: first use segmentation models to identify frames where they could identify three unique tools per frame. By identifying three tools, they could then apply the video-level labels to the given frame. In order to do this they used models [26] published from the MICCAI 2017 Robotic Instrument Segmentation Challenge [^7] https://github.com/ternaus/robot-surgery-segmentation. After applying segmentation models they then used traditional computer vision techniques to count the contours of each unique instrument, and layered additional logic (i.e., instrument contour needs to touch the frame edge) to filter out false positives. All frames where they could positively identify three unique instruments per frame then formed the basis of the training dataset.
Another unique thing in this dataset is the fact that many videos are sequential, e.g., 30-second chunks of a continuous video that last for several minutes or even dozens of minutes. Grouping videos from the same “scene” into the same fold split is critical for meaningful local validation. In order to do this, they compared the image hash of each video’s first frame and the last frame of the preceding video. If the similarity is lower than a threshold, they mark them as from the same scene and put all videos of the same scene into the same split when they split the data into 5 folds. This way, leakage between folds can be prevented. To simulate the test data, they downsampled training videos from fps=60 to fps=1. They also cropped the black empty space on the left and right sides of all the training videos. Then, they blurred out the bottom banner which may contain ground truth labels. Finally, they resized the frames into 640 W x 512 H, the same as the test videos.
After the above steps, the team trained ConvNext[27] Tiny and EfficientNet[28] B4 models, two model architectures that consistently dominate Kaggle image competitions. Standard techniques were used such as cosine learning rate schedule, mixed precision training, mixup augmentations, and weighted loss functions.
Category 2 required localization and classification of a given set of surgical instruments in each frame of a video. Only video-level labels were provided so the challenge can be tasked as a weakly supervised problem. In order to solve this problem they used a diversity of techniques taken from both supervised and unsupervised methods.
One of the key learnings from Category 1 was the importance of sampling rare classes and identifying high-confidence labels in the train set. They exploited these learnings by developing a sample methodology to get a balanced set of frames/labels from a diverse set of scenes using the models they developed for Category 1. After these frames were identified they attempted to get clean localization of the tools of interest through two methods, GradCam[29] and explicit labeling of the tools of interest. GradCam was effective at identifying the clevis of the tools of interest, however, because there was a set of rules and detection logic that needed to be applied on top of the clevis identification (i.e., tip vs clevis detection for some tools and not others), they found that applying manual labels that incorporated these rules were important. GradCam was thus used for visualization only, and manual application of labels and competition logic was used for the start of training.
After identifying clean labels, the team trained YoloV5[30] models first on a small set of manually created labels, according to the example labels on the challenge data page [^8] https://surgtoolloc.grand-challenge.org/data/, and then on weakly supervised pseudo-labels predicted from the first set of base models [31]. The workflow is illustrated in Figure 10.
Figure 10: NVIDIA: Team NVIDIA Category 2 workflow
3.5.2 Model Training
For Category 1, they trained 5 EfficientNet-B4 models (on 5 fold splits) and 5 ConvNext-tiny models and ensembled them using a trick called logit shift. The idea of the logit shift trick is that, when data is extremely imbalanced between classes as in this dataset, the minor classes’ probabilities are extremely biased towards 0. The extent of the bias can be different across models, making ensembling difficult. Logit shifting is to shift the logit of each class by an optimal constant value (i.e. the same value for the same class, across all samples), so that 0.5 probability threshold is the optimal threshold for the F1 score. This optimal logit shift is tuned on local validation, and is different for each model. When ensembling, they shift logit first on individual models then simply take the average probabilities, since they are already adjusted to be centered around 0.5.
For Category 2, the team trained several versions of YoloV5 and found YoloV5m to be the best performing. Except for the default augmentations such as mixup, mosaic, and HSV transforms, they applied more augmentations like hue saturation value and random brightness contrast to make the model generalize better, especially for the minor tools.
All the experiments for the challenge were trained on NVIDIA DGX Station A100 (with A100 GPUs) and NVIDIA DGX Station (with V100 GPUs).
3.5.3 Preliminary Performance
For Category 1, The per-class F1 validation scores of both the ConvNext and EfficientNet models are provided in Table 5, as well as the ensemble score. The final ensemble has a validation score of 0.8526 and a leaderboard score of 0.7055.
For Category 2, the team achieved a mAP validation score of 0.3614 for the first model, and 0.4132 for the retrained model, which translated to a leaderboard score of 0.3058. The validation scores for each class are shown in Table 6.
Table 5: NVIDIA: Category 1 Results
Class ConvNext Tiny EfficientNet B4 Ensemble
needle driver 0.9439 0.9565 0.9487
cadiere forceps 0.6529 0.6826 0.6907
bipolar forceps 0.9236 0.9305 0.9264
monopolar curved scissors 0.9179 0.9330 0.9189
grasping retractor 0.7082 0.6801 0.7188
prograsp forceps 0.7549 0.7946 0.7742
force bipolar 0.9216 0.8665 0.9401
vessel sealer 0.9690 0.9728 0.9805
permanent cautery hook/spatula 0.8602 0.8417 0.8489
clip applier 0.8664 0.8700 0.8439
tip-up fenestrated grasper 0.5333 0.5502 0.5222
stapler 0.9746 0.9748 0.9789
bipolar dissector 0.9787 0.9744 0.9829
suction irrigator 0.8619 0.8306 0.8619
Average F1 0.8476 0.8470 0.8526
Table 6: NVIDIA: Category 2 Results
Class Frist model Weakly supervised model
needle driver 0.5043 0.4933
cadiere forceps 0.5024 0.5111
bipolar forceps 0.3275 0.4441
monopolar curved scissors 0.3568 0.3898
grasping retractor 0.3101 0.4098
prograsp forceps 0.3619 0.4177
force bipolar 0.4341 0.4442
vessel sealer 0.4660 0.4874
permanent cautery hook 0.3866 0.4164
clip applier 0.2536 0.3520
tip-up fenestrated grasper 0.4815 0.4472
stapler 0.2262 0.2902
bipolar dissector 0.1046 0.3197
suction irrigator 0.3448 0.3617
Average mAP 0.3614 0.4132
3.6 Argonne National Laboratory - Team ANL-Surg
ANL-Surg team included Neil Getty, Fangfang Xia, Zixuan Zhao, Xiaotian Duan. This team was primarily interested in category 2, though given the available challenges, both categories were considered. ANL-Surg was motivated by the scale of the data, and the interesting need for a weakly-supervised approach. Their initial plan was to use traditional computer vision or pre-trained deep learning models to segment seen tools, and then make label assumptions about the unseen tools and retrain the model iteratively.
3.6.1 Method Description
Training data was sampled at 0.25 FPS and cropped to 880x520 to remove black borders and UI. ANL-Surg used segmentation labels from the two public competitions: MICCAI18: Robotic Scene Segmentation Sub-Challenge [19] and MICCAI17: Robotic Instrument Segmentation [25] for training initial Detectron2 segmentation models.
Their approach for tool presence classification was an ensemble of six ResNet50 models trained using the Fastai framework [32]. Five of the models were trained using cross-validation, and one was trained on all of the data. ANL-Surg averaged the softmax output and used a threshold to determine the final classification.
For tool localization, they trained a Detectron2 [33] model to segment parts of surgical tools. Then the parts were clustered to form bounding boxes for entire tools. Finally, individual tools are cropped out based on the predicted bounding box and query the Category 1 model about the class of the tool (Figure 11.
Figure 11: ANL-Surg: Top right frame is the result of the parts detection model. Bottom left is the instrument tool detection model. Bottom right is the wrist final prediction with post-processing using the Category 1 model.
3.6.2 Model Training
ANL-Surg used 5-fold cross-validation, splitting the clips based on their computed similarity scores so that clips of the same video, or very close temporally were not included in both training and validation. However, validation was still a challenge, as a good result may actually be a bad result, due to the many frames with less than 3 visible instruments.
Detectron2 models were trained on a Tesla v100s with 32gGB VRAM for 2000 iterations, a batch size of 20, learning rate of 0.02, and 4 detections per image. MaskRCNN weights were used before finetuning on MICCAI 2017 and 2018 data.
3.6.3 Preliminary Performance
Internal validation results for tool presence classification are presented in Table 7.
Table 7: ANL-Surg: Preliminary F1 Scores for 5-fold cross-validation for tool presence classification.
Fold F1
0 0.87399
1 0.91667
2 0.87041
3 0.84748
4 0.86528
3.7 Keio University - Team HVRL
Team Members: Ryo Fujii, Ryo Hachiuma, Mana Masuda, and Hideo Saito
The main motivation of the HVRL team was to familiarize themselves with weakly supervised learning in a real surgery environment. The team has experience in building large-scale tool presence detection dataset in open surgery and deeply understand how expensive and tedious to annotate the surgery video [34]. Thus, applying weakly supervised learning reduces the annotation time to a real surgery video.
3.7.1 Method Description
For Category 1, multi-label tool presence classification models were trained in a fully supervised manner. Since only video clip-wise annotation is provided, all frames in a video clip were assumed to have the same annotation as the annotation attached to the video clip. The proposed solution was the ensemble of three different models, Swin Transformer V2 [35], EfficientNetV2 [36], and ResNeXt [37], which showed the high performance to the validation data. Models were implemented based on the PyTorch Image Models  [^9] https://github.com/rwightman/pytorch-image-models [38].
For Category 2, Grad-CAM++ [39] was employed for the weakly-supervised tool localization task.
  • architecture design: For category1, the team adopted the ensemble of three different models, Swin Transformer V2 [35], EfficientNetV2 [36], and ResNeXt [37]. For Category 2, only EfficientNetV2 was employed.
    Refer to caption Refer to caption Refer to caption
    (a) Original image (b) UI cropped image (c) UI blurred image
    Figure 12: HVRL: Examples of preprocessing. The original image included the black region on the left and right sides and the tool annotation UI on the bottom. First, the original image was cropped to remove the black region. Then, UI cropping and UI blurring with a probability of 0.5 was randomly conducted.
  • data processing, reformatting, transformation, etc.: The image extracted from video clips includes a black region on the left and right sides and a tool annotation UI on the bottom. Firstly, these black regions were removed. Then, UI cropping and/or UI blurring with a probability of 0.5 were randomly conducted. These procedures prevented the model from getting the tool presence information from the UI directory. Examples are shown in Fig. 12.
  • data augmentation: several data augmentation techniques were applied for achieving robust training. Specifically, the team applied the data augmentation techniques shown in Table LABEL:tab:aug. These augmentations were applied to train all models.
  • model inputs: RGB image.
  • model outputs: For category 1, multilabel output. For category 2, pixel-wise heatmap.
  • model loss function: For category 1, binary cross-entropy loss. For Category 2, the model trained by Category 1 was used.
  • model pre-training: For category 1, The weights of Swin Transformer V2 and EfficientNetV2 were initialized with the weights trained with ImageNet [40], and that of ResNeXt was initialized with the weights trained with Instagram Datasets. For Category 2, the model trained in Category 1 was reused.
  • post-processing: For category 1, test time augmentation (TTA) was employed (Table LABEL:tab:aug). For category 2, the gradient for the last convolutional layer and the penultimate convolutional layer was computed via back-propagation. After obtaining the pixel-wise heatmap for each layer, these two heatmaps are averaged to obtain the final heatmap. Then, the heatmap is binarized with threshold $\sigma=0.5$ to obtain the binary map. Finally, the connected component algorithm is applied in order to extract multiple bounding boxes from the binarized heatmap.
  • Hyper-parameters (Table LABEL:tab:hyper_hvrl): For Category 1, Stochastic Gradient Descent was employed as the optimizer, with the momentum $0.9$, weight decay $0.0001$, and cosine annealing strategy. For the learning rate, batch size, and training epochs, different settings on three methods were used as shown in Table LABEL:tab:aug. For Category 2, the model trained in Category 1 was used, thus retraining was not conducted.
Table 8: HVRL: Data augmentation
Augmentation Range Probability
Horizontal flip 0.5
Color shift 30 0.5
HSV Shift 20 0.5
Brightness [0.1, 0.3] 0.2
Table 9: HVRL: Hyperparameters
methods learning rate batch size epochs
Swin Transformer V2 0.6 24 50
EfficientNetV2 0.57 48 30
ResNeXt 0.36 192 50
The trained model outputs a multi-class label. The Grad-CAM++ [39] module generated the class activation maps. Then, the generated class activation maps with threshold $\sigma=0.5$ generated the binary map. Finally, a connected component algorithm was applied in order to extract multiple bounding boxes from the binarized heatmap.
3.7.2 Preliminary Performance
Table 10: HVRL: Local validation results with different models.
Method F1-score
Swin Transformer V2 0.925
EfficientNetV2 0.930
ResNeXt 0.928
Model ensemble 0.930
Table 10 summarizes the results against the local validation data. Models which showed the best performance on the local validation data were chosen. Specifically on epochs 17, 25, and 23. The table shows that a single EfficientNetV2 and the ensemble of three models achieved the highest F1-score.
3.8 Muroran Institute of Technology - Team SK
Team Members: Satoshi Kondo, Satoshi Kasai, and Kousuke Hirasawa
3.8.1 Method Description
Team SK proposed a weakly-supervised surgical tool localization method with a multiplicative feature fusion network. The network was trained to learn the presence of surgical tools in the input image with supervised learning. When detecting surgical tools (task in Category 1), this method predicts the presence of surgical tools by using the trained network. When localizing surgical tools (task in Category 2), this implementation first predicts the presence of surgical tools and then checks the class activation map corresponding to the predicted surgical tool, localizing the tool by thresholding the activation map, i.e., in a weakly-supervised manner.
Figure 13 shows the entire structure of the proposed network. Multiplicative Feature Fusion Networks (MFF-Net) [41] was used as the base network. MFF-Net is designed to aggregate features at different levels, which generates the class activation map (CAM) and categorical features for classification. Then, a CAM of the size $C\times H^{\prime}\times W^{\prime}$ is obtained. Here, $C$ is the number of classes in the tool classification, and $H^{\prime}$ and $W^{\prime}$ are the height and width of the CAM, respectively. ResNet-50 is used as the backbone of the MFF-Net. The training and prediction are performed frame-by-frame base although the input data is video.
The network was then trained by using the provided dataset, which has images and labels showing the presence of tools in the images. There were 14 surgical tools and 3 tools are shown in an image at most. Based on the team’s observation that two “needle driver” tools are often shown in the same images, the model was trained not as a 14-class classification model but as a 15-class classification model, i.e., $C=15$. Since the dataset contains two “needle driver” classes in 15 classes. When only one needle driver is shown in the input image, only the first label in needle driver classes is active, i.e., give 1 as the ground truth label. When two needle drivers are shown in the input image, both labels for needle driver classes are considered active.
In inference, the probabilities of surgical tool presence were obtained from the CAMs through global averaging pooling and sigmoid activation. Tool classes whose probabilities are higher than the threshold (0.5) were selected. In this case, the selected tool classes were the results of the surgical tool detection task (Category 1). For the selected tools, this approach obtained CAM of the corresponding tool class in the order of the probability (from the highest probability to the lower ones). Then, the tool area was obtained by thresholding the CAM. The threshold was determined by using Otsu’s binarization method. After identifying the bounding box surrounding the tool area, the IoU was calculated. If the IoU value is less than the threshold, the bounding box is employed as the result of the surgical tool localization (Category 2).
Figure 13: SK: Overview of our proposed method.
3.8.2 Model Training
For the training of the network, the team constructed a dataset by sub-sampling images from videos. The total number of images for training in this set was over 150k, being split into training and validation in an 8:2 ratio.
The augmentation techniques used were horizontal flip, shift, scale, rotation, color jitter, Gaussian blur, and Gaussian noise. The augmented images were resized to $640\times 480$ pixels. The employed optimization method was Adam, and the initial learning rate was set to $1.0\times 10^{-5}$ changing at every epoch with cosine annealing. The cross-entropy loss was used as the loss function. The batch size was set to 64 and the number of epochs to 30. The model taking the lowest loss value for the validation dataset was selected as the final model. Five models were trained by using 5-fold cross-validation, and CAMs obtained by five models were averaged in inference. An NVIDIA RTX3090 GPU for the training.
3.8.3 Preliminary Performance
When using the evaluation system provided by the challenge organizers, the F1 score was 0.816 for Category 1 (surgical tool detection). For Category 2 (surgical tool localization), mAP was 0.03.
3.9 Team VANDY-VISE
Team Members: Xing Yao, Ange Lou, Hao Yang, Jintong Han, Jack Noble, Jie Ying Wu
The goal was to take full advantage of the auto-correlation and cross-correlation between various frame-to-frame and video-to-video to solve this weakly supervised recognition problem. Therefore, an attention-based learning framework was selected to capture and detect causal features that were directly related to surgical tool detection. No public dataset apart from the challenge dataset was used in this work.
3.9.1 Method Description
Figure 14: VANDY-VISE: Architecture of Query2Label [42]
The VANDY-VISE team leveraged the same framework as Query2Label[42] to do multi-label classification and localization (shown in Figure 14). The input image $x$ was first sent through a ResNet backbone [43] to extract the spatial features. The extracted spatial features were then sent into the transformer along with each of the label embeddings. The transformer decoders generated attention maps by comparing label embeddings with spatial features, then pooling the spatial features by linearly combining these spatial features based on the attention map. After obtaining spatial features of the input image, each label embedding was sent to the transformer decoder to query and pool the desired feature. Finally, those features went through a linear projection layer to compute the logits for each class.
For a given image $x\in R^{H\times{W}\times{3}}$, a ResNet backbone was used to extract spatial features $f_{0}\in R^{H_{0}\times{W_{0}}\times{d_{0}}}$ from the input image. Here, $H\times{W}$ and $H_{0}\times{W_{0}}$ represented the height and width of the input image and extracted features respectively, while $d_{0}$ denoted the dimension of features. The spatial features were then projected from $HW\times d_{0}$ to $HW\times d$ by using a linear projection layer, where $d$ was the desired feature dimension.
The class embedding was applied as a query that had dimension $Q^{K\times d}$. Here $K$ represented the number of classes. Then the cross-attention was calculated between the spatial features and queries to pool the class-related features by using multi-layer transformer decoders. In the implementation carried out by the Vanderbilt University team, the standard transformer decoder comprised cross-attention, self-attention, and a position-wise feedforward network. The transformer decoder, $i$, updated the queries $Q_{i-1}$ as following:
$\operatorname{self-atten}:Q_{i}^{(1)}=\operatorname{MultiHead}(\tilde{Q}_{i-1},\tilde{Q}_{i-1},Q_{i-1})$
$\operatorname{cross-atten}:Q_{i}^{(2)}=\operatorname{MultiHead}(\tilde{Q}_{i-1}^{(1)},\tilde{f},f)$
$\operatorname{FFN}:Q_{i}=\operatorname{FFN}(Q_{i}^{(2))}$
where the hat represented variables with position encoding.
The query feature $Q_{i}$ was sent to a linear projection layer to a logit value for each class.
$p_{k}=\operatorname{Sigmoid}(W_{k}^{T}Q_{i,k}+b_{k})$
in the right-hand-side of the function, $W_{k}$ and $b_{k}$ were parameters of the linear projection layer. The left-hand side, $p_{k}$, was the predicted probability of each class.
In the implementation, a simplified asymmetric loss [44] was applied as shown below:
$L={1\over K}\sum_{k=1}^{K}\left\{\frac{(1-p_{k})^{\gamma+}log(p_{k}),y_{k}=1}{(p_{k})^{\gamma-}log(1-p_{k}),y_{k}=0}\right\}$
here, $y_{k}$ was a binary label for $k$ classes, which corresponds to the $\gamma$ value to choose. By default, they are set to $\gamma+=0$, $\gamma-=1$. It was noteworthy that countless noisy labels existed in the training annotations, and it was very time-consuming to remove them manually. Most of the label noise came from pseudo-labels that were not present in the videos. Based on this prior knowledge, during the training phase of the classification model, for videos with more than two labels, the team removed the largest loss among all existing labels. Compared with the original loss, this method improved the best validation mAP from 0.953 to 0.966.
In this work, the VANDY-VISE team post-processed the cross-attention weights created by the transformer decoder to generate bounding boxes for each predicted class. The post-processing included sparse matrix processing and bounding box generation based on OpenCV [45]. During the attention visualization process, the cross-attention weights matrix $M$, produced by the decoder, was initially reshaped from dimensions $B\times C\times HW$ to $B\times C\times H\times W$. This transformation enabled the representation of cross-attention between each pixel and its respective class. Then, according to the classification results of the model, the classes in the normalized attention map were filtered, and only the classes with predicted True were retained. Because tools of different categories shared similar morphological features, the attention maps of different categories still had similarities. To further distinguish the critical vector of each class, for different positions of the same class, this method selectively retained one or several positions with the highest energy; for the exact position of different classes, it only retained the position corresponding to the class with the highest energy. Finally, the team binarized the sparse attention maps and used cv2.findContours and cv2.boundingRect [45] functions to generate corresponding bounding boxes.
3.9.2 Model Training
To make the problem more tractable, one frame from each video clip was randomly selected to build a relatively smaller dataset, which in total contained 24695 images. From this whole set, 3000 images were randomly partitioned as the validation set, and the rest 21695 frames were assigned to the training set. Blank areas in each frame were removed. Furthermore, AdamW was chosen as the optimizer with a learning rate of 1e-4. The training batch size was 32. All experiments were based on Pytorch and trained with a single RTX A5000 GPU.
3.9.3 Preliminary Performance
In multi-label classification tasks, the loss decreased steadily during the training phase, and the Vanderbilt University team achieved 0.966 mAP and 0.955 F1 scores on their validation set and a 0.6202 F1 score on the given test set. For the localization task, their algorithm (multi-head attention weights map) could localize all surgical tools that appear in the given validation video.
3.10 University Medical Center Hambrug-Eppendorf - Team UKE
Team Members: Maximilian Nielsen, Samuel Schüttler, Thilo Sentker, Hümeyra Husseini, Ivo Baltruschat, Rüdiger Schmitz, René Werner
3.10.1 Method Description
Recently, self-supervised deep learning gained attention by setting new benchmarks for different image classification tasks, showing potential for further investigation in related domains [46, 47]. Different to supervised methods, self-supervision does not directly produce a meaningful output (e.g., a classification result or segmentation mask), but solely a representation, i.e. a feature-vector, of the input. To find an expressive representation of the input, such methods predominately rely on contrastive learning, which is characterized by forcing the model to output similar representations for different augmented versions of the same input. Advantages compared to supervised learning approaches are a relative strong performance with a fraction of labeled data and a high degree of generalizability. Therefore, a self-supervised approach for both tasks of the SurgToolLoc challenge was employed. More specifically, deep image representations are extracted by self-supervised vision transformers (DINO). Subsequently, downstream classification is performed by a machine-learning classifier that is fitted on the model output of a smaller subset of labeled data. For tool localization, the individual outputs of the trained attention heads to generate bounding boxes were exploited.
For both tasks, the starting point was to extract feature vectors exploiting a self-supervised learning approach based on DINO (cf. figure 15 for a schematic illustration). Prior to training, frames were resized to $308\times 208$ px and frame intensities normalized to ($-1,1$). A DINO model, pre-trained on the ImageNet data, was trained on the given pre-processed training data with slightly adapted default DINO-parameters.
Several image augmentation techniques were utilized to improve the robustness of the model. The DINO augmentation techniques included random cropping to simulate different perspectives, solarization to introduce variations in brightness, flipping to incorporate left-right reversal, Gaussian blurring to mimic real-world blurs and color jittering to introduce variations in color.
A temporal global crop was applied to consider the temporal components of the input data. This approach involves randomly selecting a global crop from a frame that is in temporal correspondence with the input frame, with a tolerance of ±$\pm 5$ s. This approach ensures temporal consistency between the input frame and the selected global crop, i.e. the temporal structure of the data is preserved which further diversifies the input data.
The loss function used was the DINO loss, which is a combination of temperature-weighted cross-entropy and feature normalization. The temperature-weighted cross-entropy term ensures a sharp feature distribution in the latent space, while the feature normalization term stabilized the training process and improved generalization by enforcing uniformity.
For task 1, a machine-learning classifier (logistic regression) was fitted on the DINO model output of a smaller subset of labeled data to learn the relationship between extracted feature vectors and corresponding training labels. Hyper-parameter optimization, feature selection and decision threshold optimization using a differential evolution algorithm were performed.
For task 2, a linear combination between multi-head attentions, RGB-values and minimum RGB-value to gray-value-distances was optimized to match the EndoVis 2017 instrument segmentations (shaft, tip, clevis) [25]. Subsequently, bounding boxes for the three largest connected components were extracted from the resulting segmentations. For each crop, a pseudo-label was created from possible ground-truth values and similar other crops (DINO latent space). The tool classifier was trained as in task 1.
Figure 15: UKE: DINO-based self-distillation. Global and local crops are extracted from the input frame, individually augmented and fed into the teacher and student model, respectively. Note that a temporal global crop was used, i.e. the global crop is randomly extracted out of one frame that is in temporal correspondence to the input frame ($\pm$5 s) to achieve temporal consistency. Further, the applied softmax function has a temperature parameter to control the sharpness of the output probabilities. After training, a machine-learning classifier is fitted on the DINO model output of a smaller subset of labeled data. For more details cf. the corresponding DINO-publication [46].
3.10.2 Model Training
The model was trained for a total of 30,000 epochs using a batch size of 224 on a GPU server equipped with an A40 GPU, 1TB of RAM, and a CPU with 64 cores. Model training took about 35 hours.
To ensure a deterministic output and a fair comparison between different runs, a fixed random seed, i.e. sum of int32 interpreted sha256-hash of string “STL-Challenge”, was used throughout model training. Thus, consistent results for each run were obtained and it was possible to compare the performance of different models in a controlled environment.
3.10.3 Preliminary Performance
Applying the proposed algorithm on the self-assessed validation split containing 30% of the provided video data yielded a F1-score for task 1 of 0.82. For task 2, validation was not possible as ground truth values were not available.
3.11 Furtwangen University - Team ITeM
Team Members: Tamer Abdulbaki Alshirbaji, Nour Aldeen Jalal, Herag Arabian, Ning Ding, and Knut Moeller
The ITeM team is composed of four PhD students who are currently working under the guidance of Prof. Knut Moeller at the Institute of Technical Medicine, Furtwangen University, Germany. ITeM team members are involved with research projects in collaboration with industrial and clinical partners. These projects aim to develop and integrate smart technologies into the current setup inside the operating rooms (OR). Analyzing surgical videos using deep learning approaches, particularly surgical workflow recognition, represents a core topic that has been addressed by the researchers at ITeM. Spatio-temporal deep learning approaches for surgical tool classification and phase recognition have been developed. Additionally, data fusion approaches have also been investigated on data from surgery and anesthesiology. Therefore, participation in the SurgToolLoc challenge represents a great opportunity for the ITeM team to evaluate developed approaches on a new challenging dataset and to be in contact with other highly qualified teams from around the world.
3.11.1 Method Description
The SurgToolLoc challenge addresses supervised tool classification and weakly-supervised tool localization tasks in laparoscopic videos. In both tasks, only binary labels for tool presence can be used to train the proposed model. Similar to other public datasets (e.g., Cholec80 [20]), surgical tool presence is imbalanced, which will cause a bias on the trained model to high-representative tools. Additionally, tool classification is a multi-object classification task, where up to 4 tools may appear in the laparoscopic image. Moreover, the model will be tested with images from video clips that are not included in the training data.
A deep learning pipeline that addresses all the above-mentioned challenges was developed. The model consisted of the ResNet-50 architecture as a backbone model since it showed better performance over other CNN architecture for the tool classification task [48]. For weakly-supervised tool localization, the algorithm was built upon the work of Durand et al. [49] and the follow-up approaches by Vardazarya et al. [50] and Nwoye et al. [51], but introduced novel modifications. First, attention modules were adopted to the base CNN model to help generate better-refined features. Second, features from multiple stages (i.e., from lower and top layers) were fused for a better representation of laparoscopic images’ contents. Third, a novel spatial pooling was applied to obtain tool presence confidence from the localization maps. Finally, a new loss function was proposed to model possible tool combinations of each video frame (Figure 16 shows the architecture of the proposed model).
Figure 16: ITeM: The complete pipeline of the proposed model for tool presence detection and localization. SE blocks represent the squeeze-and-excitation attention modules; the Conv2D layer is the multi-map localization layer; FC-tool and FC-Combination layers are the fully-connected layers employed to obtain confidence values of the tool presence and tool-combination, respectively. N is the number of tool classes defined in the dataset (N=14).
  • architecture design:
    • base model
      The proposed architecture was built on the base CNN model ResNet50 [43]. This model showed competitive performance for surgical tool classification in laparoscopic images [48]. ResNet50 consists of five convolutional blocks and on top a global average pooling and a fully-connected layer. The shortcut connections used in the model lead to more efficient optimization and higher accuracy [43]. The model input was increased to 375×300 instead of 224×224. Additionally, the stride of convolutional layers was set to 1x1 in the last two blocks. Due to these modifications, higher spatial resolution is preserved at deeper layers.
    • squeeze-and-excitation module
      Squeeze-and-Excitation (SE) [52] attention modules were integrated into the convolutional architecture to boost the model performance. Four SE modules were added after the fourth, seventh, thirteenth, and sixteenth convolutional blocks. The reduction ratio was set to 16 in all SE modules.
    • multi-stage feature fusion
      In conventional methods, the CNN model performs the classification based on information learned at the last layer. In fact, features of deeper layers encode semantic information about classes. On the other hand, features of shallow and intermediate layers encode some generic information. Feeding such information for instance to a classification layer can support the classification decision [53]. Thus, in this approach, features at the output of the second, third, and fourth attention modules were combined forming a feature bundle. The combined features were passed to a batch-normalization layer.
    • Conv2D layer
      A convolutional layer was used to transfer the multi-stage features to a localization map for each class. The convolutional layers had a kernel size of 3x3. The number of filters was set to MxN, where M is the number of maps for each class and N is the number of tool classes (N=14). M was set to four, thus, four localization maps were generated for each surgical tool. In these convolutional layers, the stride was set to 1x1 to preserve the spatial resolution.
    • pooling operations
      To generate one localization map and presence confidence for each tool, tool-wise pooling, and spatial pooling operations were applied consecutively. The max operation was applied across the 4 localization maps of every tool. The input and output of this operation are $w{\times}h{\times}4N$ and $w{\times}h{\times}N$, respectively, where $w$ and $h$ represent the spatial dimension of the obtained feature maps, while $N$ is the number of tool classes in the dataset.
      Special spatial pooling (equations 3.11.1, 3.11.1) was then applied to transform the map of each tool into two features. These two features represent the positive evidence ratio (PER) and the negative evidence ratio (NER), and they were provided into a fully-connected layer with a Sigmoid activation to obtain tool-wise confidence values.
      $PER=\frac{\sum_{i}^{h}\sum_{i}^{w}x_{i,j}}{h.w}\end{aligned}\qquad\Longleftrightarrow\qquad\begin{aligned} x_{i,j}=1&\text{ for }x>0.9\\ x_{i,j}=0,&\text{ Otherwise}\end{aligned}$
      $NER=\frac{\sum_{i}^{h}\sum_{i}^{w}x_{i,j}}{h.w}\end{aligned}\qquad\Longleftrightarrow\qquad\begin{aligned} x_{i,j}=1&\text{ for }x<0.1\\ x_{i,j}=0,&\text{ Otherwise}\end{aligned}$
      where $h$ and $w$ are the spatital dimension of the localization maps, $x_{i,j}$ is the pixel value at $i$ and $j$ position.
    • fully-connected (FC) layers
      Two fully-connected (FC) layers were added on top of the spatial pooling layer. The first FC layer (FC-tool) performs the tool presence detection and contains 14 nodes since 14 tools are defined in the dataset. The confidences obtained from the FC-tool were then provided into the second FC layer, termed FC-Combination), which was added to model the possible tool combinations in the dataset. Therefore, tool-combination classes were generated for all images prior to training the model. First, all possible tool combinations were extracted from the provided tool classes with their corresponding distributions. Then, 14 tool combination classes were generated. The first 13 classes represent the combinations of the highest thirteen distributions and the last class represents the remaining combinations.
  • data pre-processing: The video clips provided had a frame rate of 60 frames per second (FPS). Training images were extracted at 1 fps. Since some clips contained black frames, these frames were identified and excluded. The label provided for every clip was assigned to its corresponding frames.
    The image pre-processing included cropping out the black sides of the images, blurring the user interface (UI) information appearing at the bottom of the images, resizing the image into the model input (375×300×3), and image normalization.
  • model input: The input of the model is an image of size 375×300×3.
  • model outputs: The outputs of the model are confidences of the 13 tool categories defined in the dataset and localization maps of detected tools.
  • model loss function: Two loss functions were employed in this study. The binary cross-entropy function was used to compute the loss of each tool class, as in Eq. 3.11.1. To reduce the imbalanced data effect, loss-sensitive learning approach [54] was applied. Here, the loss of each tool was multiplied by a weighting factor that was estimated according to the tool-appearance distribution in the training data. Additionally, the Softmax cross-entropy loss was utilized to calculate the classification error for the tool combination (Eq. 3.11.1).
    $Tool_{loss}=\frac{-1}{B}\sum_{n=1}^{B}\sum_{t=1}^{T}w_{t}[l_{t}^{n}\log{(C_{t}^{n})}+(1-l_{t}^{n})\log{(1-C_{t}^{n}})]$
    where $Tool_{loss}$ is the total loss of all tools, $B$ is the batch size, $T$ is the number of tools in the dataset, $w_{t}$ represents the loss-weight calculated for every surgical tool, $l_{t}^{n}=[0,1]$ is the tool presence ground truth, and $C_{t}^{n}$ is the tool presence confidence obtained from the spatial pooling operation.
    $loss_{comb.}=\frac{-1}{B}\sum_{n=1}^{B}\sum_{c=1}^{C}G_{c}^{n}\log{\sigma(C_{c}^{n})}$
    where $loss_{comb.}$ is the total loss of tool-combination category, $B$ is the batch size, $C$ is the number of defined tool-combination in the dataset, $G_{c}^{n}$ is the ground truth of image $n$, $C_{c}^{n}$ is the output of the FC-Combination layer, and $\sigma$ represents the softmax activation function.
  • model pre-training: The base ResNet-50 model was pre-trained with the Cholec80 [20] dataset. The weights of the newly added layers were randomly initialized.
  • post-processing: post-processing step was required to extract bounding boxes from the localization maps generated after applying the tool-wise pooling operation. Localization maps of the tools that were detected to be present in the image (confidence >0.5) were extracted. A Gaussian filter with sigma = 1.0 was applied to smooth boundaries in the localization maps. Then, a binary mask was generated by thresholding the localization map with a threshold of 0.5. Bounding boxes were finally generated on top of generated masks. For the needle driver tool, up to two bounding boxes were considered.
3.11.2 Model Training
The Adam optimizer with cyclical learning rate techniques was utilized. The cyclical learning rate boundaries were chosen as 0.005 and 0.001, and a step-size of $4\times iteration-per-epoch$ was chosen. The model was trained with a batch size of 50 images for 50 epochs. Images were shuffled for every epoch. All implementations and experiments were accomplished in Keras framework with Anaconda virtual environment platform and run on a desktop with an AMD Ryzen Threadripper PRO 3955WX 16-Cores 3.90 GHz, 512 GB RAM, and 64-bit Windows 10 Operating System with an NVIDIA RTX A6000 graphics processing unit (GPU).
3.11.3 Preliminary Performance
To evaluate the trained models prior to submission, a small, denoised-label dataset was created by manually revising the provided labels. Figure 17 shows the performance of the proposed model for surgical tool classification on the validation set. Additionally, the results of different models on the preliminary-test data are presented in Table 11.
The experimental results of the proposed model showed high performance on both the validation and preliminary test data. However, an extensive ablation study is still required to show the advantage of every component of the model. Moreover, additional work should be done to denoise the tool labels provided.
Figure 17: ITeM: Model performance on the validation data.
Table 11: ITeM: The results of the different models obtained from the preliminary test
Model F1-Score (%)
Base model 51.76
Base model+Conv2d+Pooling 76.23
Base model+Conv2d+Pooling+SE 77.51
Final Model 84.35
3.12 Shun Hing Institute of Advanced Engineering - Team Medical Mechatronics (MM)
Team Members: An Wang, Mengya Xu, Mobarakol Islam, Long Bai, Winnie Pang, Hongliang Ren
3.12.1 Method Description
Multi-label tool classification from surgical videos is challenging but demanding to increase the autonomy of computer-assisted interventions. The method from Medical Mechatronics (MM) utilized the hybrid ViT [55] model (R50-ViT-B_16)  [^10] https://github.com/mobarakol/Hybrid_ViT to complete the classification task. Further, the team considered solving the noise and imbalance issues of the dataset by intentionally splitting the dataset into training and validation and with an unbalanced dataset sampler during training. They also applied augmentations like Cutout [56], and RandAugment [57], to increase generalization.
There are always dataset issues like noisy labels, imbalanced datasets, and domain shifts in the practical dataset. It is important to consider proper data cleaning and manipulation before feeding the data to the deep learning models. The entire training dataset consists of over 20 thousand videos. There are 14 types of tools in these videos. The frequency of their appearance is quite different, making the dataset imbalanced. The first frame from each valid video (except for 9 corrupted videos) was extracted to represent each video clip to reduce redundancy and save memory. The corresponding label of each frame is a one-dimensional array with a length of 14. The element of the respective class was set to 1 if the tool exists and 0 otherwise. The dataset was not randomly split into training and validation sets. Instead, the team analyzed the dataset and ensured all tools existed in the training and validation periods. The number of training samples was about 80% of the entire dataset, and the number of each class of samples for training was also around 80%.
The training images were resized to 224X224, and two advanced augmentation techniques, i.e., Cutout and RandAugment, were incorporated as data augmentation to avoid overfitting. The PyTorch-based multilabel balanced sampler  [^11] https://github.com/issamemari/pytorch-multilabel-balanced-sampler was adopted to sample the samples during training. Asymmetric Loss (ASL) [58] was used as the multi-label classification loss function, which performs differently on positive and negative samples by automatically reducing the weight and hard limit of easy negative samples, and eliminating samples that could be mislabeled. Figure 18 describes their model architecture.
Figure 18: MM: R50+ViT-B_16 hybrid model
3.12.2 Model Training
During training, SGD was used as the optimizer, and a cosine scheduler was applied to the learning rate (max learning rate was 3e-2) for 1000 warm-up steps. The total number of training steps was 10000. The hybrid ViT model was trained on one RTX3090 GPU with a batch size of 128, taking around 3.5 hours to finish the training. The balanced multi-label sampler in the dataloader helped achieve training-time class balancing. The evaluation interval was set to 50 steps, and the final model was acquired from the step with the highest validation MAP.
Figure 19: MM: The training process.
3.12.3 Preliminary Performance
An mAP of 0.86 was achieved on the validation dataset. Figure 19 shows the training process.
3.13 Bin Zayed University of Artificial Intelligence - Team BioMedIA
[…]
The team was focused on the task of tool classification but they targeted the localization task as well considering the extent to which the task of localization could help the classification problem. The authors experimented with simple architectures like ResNet 50 first and planned to build on these simple architectures to increase the performance of the model. In the end, the team used a combination of two architectures: ConvNext and Yolov5 which helped them improve the performance metrics.
3.13.1 Method Description
The authors developed a method that localized different surgical instruments in surgical videos with a two-stage model, a detection foundation model based on YOLOv5 [59] to derive the bounding boxes and a classifier based on the ConvNext [27] architecture.
  • Architecture Design: The authors proposed a two-stage framework by combining two standard architectures YOLOv5 and ConvNext. Yolov5 was used for the task of object detection. ConvNext, on the other hand, was used for image classification. The training was organized in a way that would reduce the effect of noisy labels.
  • Data Preprocessing: According to the authors, the dataset posed many challenges due to the following reasons:
    • Dataset Imbalance: As per the data exploration done by the authors, the imbalance in the dataset was huge. The videos contained some instruments which appeared thousands of times as compared to the others which appeared only tens of times as shown in Figure 20. The authors balanced the dataset by sampling videos with different frame rates taking into account the types of tools present in each video. They also optimized (minimized) the standard deviation of the frequency of each tool in the resulting dataset.
    • Presence of banners: The banners showed the tools used in the surgery, and as proposed by the authors, the removal of the banners was crucial as it could have caused the model to learn the labels from the banners and not actually identify the instruments for their proper classification.
    • Noisy Labels: The authors, upon careful inspection of the videos in the dataset, discovered that some tools might never appear in the video even though the label suggested otherwise, which made the labels noisy and the training challenging.
    Figure 20: BioMedIA: Frequency of each tool across the whole dataset. The graph illustrates the frequent appearance of some tools compared to the other tools. This creates a huge imbalance in the dataset.
  • Data Augmentation: Due to the nature of the dataset, where numerous frames were similar, the authors proposed different transformations that could help in training. Different types of augmentations were used for the detection model and the classifier model. For detection, the authors used mosaic augmentation, mixup augmentation, and random perspective transformations. On the other hand, for the classifier, the authors used random image flipping, rotation, and color jittering.
  • Model Input(s): As an end-to-end model, the proposed framework was designed to have a video frame (RGB image) as an input. The frames were obtained by processing the video data provided.
  • Model Output(s): For each of the input frames, the framework independently outputs a set of labels for surgical tools present in the respective frame.
  • Model Loss Function: The authors used a combination of box regression loss and objectness loss in order to train the localization model and a focal loss as the criterion for the classifier in order to partially mitigate the impact of the huge class imbalance in the dataset.
  • Model Pretraining: The authors maintained that pretraining the model on different datasets that had similar characteristics as the provided data would, to some extent, boost the performance of the models. They trained ResNet50 and ConvNext on Cholec 80 [20] and M2CAI [21]. Cholec 80 dataset consists of 80 cholecystectomy surgeries. The labels present are primarily phase and tool presence annotations, where the ground truth for phase annotations is shown in Figure 21 and has been confirmed by a surgeon. On the other hand, the m2cai16-tool-locations dataset contains 15 videos with tool annotations for seven surgical instruments, as shown in Figure 22. The labels are given per frame for 7 classes for each video.
    Figure 21: BioMedIA: Surgical instruments present in Cholec80 Dataset. The tools are similar to the tools present in the challenge dataset.
    Figure 22: BioMedIA: Surgical instruments present in M2CAI Dataset. Some of the tools present are similar to the ones present in the given dataset.
For the task of classification, out of various CNN architectures, ConvNext-Small provided the best results. However, the authors made the following assumptions about the data:
  1. 1.
    They considered and filtered frames where all three instruments were present. They assumed that these frames were sufficient for the task of classification. They also assume that the distribution of surgical tools in these frames would be approximately the same as that in the actual dataset.
  2. 2.
    In the vast majority of frames, the order in which the tools are listed on the label of a clip corresponds to the counter-clockwise order of their positions, starting from the upper left corner and ending with the upper right corner.
Training algorithm: The authors used the localization model Yolov5 pre-trained on a combination of three datasets: EndoVis’15 Instrument sub-challenge Dataset, CholecSeg8k [60] and Endoscopic instrument segmentation with crowdsourced data challenge dataset [61] to detect the surgical tools in a frame with bounding boxes. The bounding boxes were saved as separate images alongside their respective image-level ID number derived from their position relative to the center of the image in an anti-clockwise fashion. The bounding boxes are then assigned the labels from the corresponding video clip (from left to right, depending on the id number of the bounding box). The outputs obtained were a set of predictions for the tools in the frame (Figure 23.
Figure 23: BioMedIA: Architecture of the two-tier model used for classification. Images are passed through Yolov5 for generating bounding boxes around the tools. Then they are passed to ConvNext for classification.
3.13.2 Model Training
To train the detection model, the authors used a batch size of 128, a learning rate of 0.001, and a total number of epochs of 300. The classifier was using a batch size of 128, a learning rate of 0.0001, and the number of epochs was set to 5 due to the fast convergence of the network.
Given the size of the dataset, the authors used a cluster to run their experiments on a GPU with 2x AMD EPYC 7742 CPUs (128 cores total with 256 threads), 256GB RAM, 4x Nvidia A100 SXM 40G GPUs.
3.13.3 Preliminary Performance
The performance of the model during training was quite high with models like ResNet-50 and ConvNext. When pre-trained on the Cholec-80 dataset, ResNet produced a training F1 score of 0.989, while for ConvNext it was 0.985. Moreover, both of them produced around the same F1 scores when trained on the M2CAI dataset. The authors believe that the unrealistically high-performance metrics were due to the similarity across the frames in different surgeries and the usage of the same set of tools for different surgeries with the same background. The validation performance for ConvNext-Small and the proposed two-tier model is shown in Table 12. However, the testing performance for the proposed architecture was around 0.552.
Table 12: BioMedIA: Comparison of different models on the performance metrics. Models were pre-trained on Cholec 80 and M2CAI.
Model F1-Score Accuracy
ResNet-50 $0.20$ $0.10$
ConvNext-S 0.83 $0.39$
Yolov5+ConvNext-S $0.75$ 0.50
3.14 University of Tokyo - Team WhiteBox
Team Members: Zhenqiang Li, Yoichi Sato
The authors introduced a network architecture for multi-label surgical tool classification from endoscopic videos. To address the challenge of subtle variance in tool appearance, the network utilizes feature maps from an intermediate layer of ResNet, which retain a higher level of spatial detail. Additionally, the long-tailed class distribution of tools is addressed by employing the asymmetric focal loss in place of the commonly used binary cross-entropy loss.
3.14.1 Method Description
The proposed network architecture comprises a feature map extraction module based on the ResNet-18 architecture and a feature post-processing module for multi-label tool classification. A significant challenge in this task is the subtle variance in the appearance of various tools. To address this issue, the network utilizes the feature maps generated by an intermediate layer of the ResNet architecture, as opposed to the final feature obtained after the 2D Average Pooling layer of ResNet. This practice is motivated by the observation that the feature maps obtained at this intermediate stage retain a higher level of spatial detail, thereby enabling the detection of tools with fine variance in appearance. Another critical challenge in this task is the long-tailed class distribution of tools, where the number of samples for the most frequent tool class in the training set is 1149 times that of the least frequent one. To address this imbalance in the training data, the network utilizes the Asymmetric Focal Loss (ASL) [58] in place of the commonly used binary cross-entropy loss as the training loss.
The overall framework for classifying tools’ presence in each frame is illustrated in Figure 24. The network utilizes the cascaded first four convolutional blocks of ResNet to extract feature maps for each frame, which have a higher spatial resolution and retain more details in the frame compared to those extracted from the fifth convolutional block or the final average pooling of ResNet. A post-processing module, consisting of an average pool layer (AvgPool) and a Multi-Layer Perceptron (MLP), is utilized to aggregate features in spatial dimensions and conduct feature transformation. The presence probabilities for 14 tools are then predicted on each frame by feeding the post-processed feature into a fully-connected layer (FC) and a Sigmoid activation layer.
Figure 24: WhiteBox: The tool presence classification framework used by WhiteBox team.
The network is trained via asymmetric focal loss (ASL) to overcome the long-tailed class distributions in the training dataset. ASL adapts focal loss for the multi-label classification task by decoupling the focusing levels of the positive and negative samples. Given $K$ labels, the network predicts one probability per label, $p_{k}$, whose ground truth is denoted by $y_{k}$. The total classification loss $L_{total}$ is obtained by aggregating the binary loss from $K$ labels as $L_{total}=\sum^{K}_{k=1}-y_{k}L_{+}(p_{k})-(1-y_{k})L_{-}(p_{k})$. The positive and negative loss parts can be formed as follows by following the focal loss:
$\begin{cases}L_{+}=(1-p)^{\gamma}\log{(p)},\\ L_{-}=p^{\gamma}\log{(1-p)},\end{cases}$
where p is one predicted probability. ASL incorporates two key modifications to the loss in order to improve the multi-label training effect. First, ASL assigns distinct weights, denoted as $\gamma_{+}$ and $\gamma_{-}$, to the positive and negative loss parts, respectively. By emphasizing the contribution of positive samples through setting $\gamma_{-}>\gamma_{+}$, ASL ensures that the network is able to learn meaningful features from positive but infrequent samples. Additionally, in order to reduce the ratio of very easy negative samples on the overall loss, ASL fully discards negative samples when their probability is very low by performing the hard thresholding as $p_{m}=\max(p-m,0)$. Here $m\geq 0$ is a tunable hyperparameter for the probability margin. In general, ASL is defined as follows:
$ASL=\begin{cases}L_{+}=(1-p)^{\gamma_{+}}\log{(p)}\\ L_{-}=(p_{m})^{\gamma_{-}}\log{(1-p_{m})}.\end{cases}$
3.14.2 Implementation Details
During the training phase of the network, the video frames were resized to a height of 270 and a width of 480. Subsequently, the upper-middle regions (shaped in 240$\times$336) of the resized frames were cropped as input to the network. This cropping process removed the black regions present on both sides and the text labels located at the bottom of the frames. Each training sample for one video was formed by splitting it into 32 segments and randomly sampling one frame from each segment. At the network test phase, the continuous regions with pixel values less than 1, which represented the black background, were first identified and removed from the frames. The remained foreground was then resized to 270$\times$336, from which the upper 240$\times$336 regions were cropped as input to the network. All frames of a video were taken as input for making predictions at the test phase. The input RGB frames were normalized using the means of (0.432, 0.395, 0.376) and the standard deviations of (0.228, 0.221, 0.217). The network outputted the existence probabilities of all 14 tools, from which the first 3 classes having probabilities larger than 0.5 were selected as the predicted tool labels.
The framework employed the first four convolutional blocks of a ResNet-18 model pre-trained on ImageNet as the feature extractor. To preserve the high spatial resolution of the output feature maps, the max pooling layer following the first convolutional block is eliminated. The resulting feature maps have a spatial shape of 15$\times$21 and 256 channels for a 336$\times$240 frame. The MLP for post-processing preserves the feature channel number at 256. These features are finally fed through a linear transformation network and a Sigmoid activation layer to predict frame-wise tool existence probabilities.
The parameters of all components were trained in an end-to-end manner using a Stochastic Gradient Descent (SGD) optimization algorithm for a total of 60 epochs. The learning rate was initially established at 0.01 and decreased by a factor of 10 after the 40th epoch. The training process was conducted using an Nvidia A100-SXM4-40GB GPU, with a batch size of 18. Additionally, the weights for the positive and negative samples in the Asymmetric Focal Loss (ASL) were set as $\gamma_{+}=1.0$ and $\gamma_{-}=4.0$, respectively. Furthermore, a probability margin of $m=0.05$ was employed to constrain the easy negative samples.
3.15 University of Strasbourg - Team CAMMA
Team Members: Chinedu Nwoye, Luca Sestini, and Nicolas Padoy.
3.15.1 Method Description
To detect surgical tools from videos acquired from Da Vinci robotic system, a spatial attention network (SANet) was proposed. This was motivated by the characteristics and constraints of the provided dataset: (1) the videos were annotated with binary presence labels from the Da Vinci user interface (UI) - in this case, a set of tool labels was fixed for a video with a high level of inconsistency when considered at a frame level, (2) utilizing the UI information was not allowed for the SurgToolLoc challenge and would also be blurred for the test videos, and (3) the test labels would be frame-based, more accurate, and thus, not a fair representation of the provided training videos and labels. Taking these limitations into consideration, the SANet model was designed to discover more useful features that could represent the tools’ presence and yet with some level of tolerance to the noisy ground truths.
The model and its implementation details are summarized as follows:
  • Architecture design: The SANet consists of 3 modules as shown in Figure 25: (1) a ResNet-18 [43] backbone to extract spatial features from the input images, (2) a spatial attention module, consisting of 8 layers of a self-attention unit [62], to highlight the discriminating regions of the input features for the tools with each feature channel constrained by designed to focus on a distinct tool class, and (3) a global pooling-based classifier [63] to transform the attention features to a vector of class-wise probabilities for the 14 tool classes considered in the dataset.
  • Data processing: The images were cropped to eliminate UI sections before being resized to a $360\times 640$ spatial dimension because the video frames contain a sort of tool information from the Da Vinci UI that would not be available during testing.
  • Data augmentation: The input images were augmented using color jittering, flipping, random rotation, and random adjustment of brightness/contrast.
  • Model input: During training, the model received pre-processed and augmented inputs, while during inference, it received solely pre-processed inputs.
  • Model outputs: The output was a vector representing the log probability of each tool category appearing in the image. The output attention maps also included location-based activations for tool instances, which helped localize the tools in a weakly-supervised scenario.
  • Model loss function: The model was trained using weighted sigmoid cross entropy loss. The class weights were computed by median-weighted inverse frequency given in [63].
  • Model pre-training: The ResNet-18 backbone was pretrained on ImageNet [22].
  • Post-processing: The output vector of the model was converted to binary labels at a threshold of 0.5 and the positive class identities were mapped to the tool labels.
Figure 25: CAMMA: Architecture of the spatial attention network (SANet) for surgical tool presence detection
Using its ResNet-18 backbone, the SANet model extracts spatial features from an input image. The spatial self-attention module highlights the feature regions that belong to the tools while suppressing those that belong to the background. The classifier pools the maximal activations of the feature channels into a vector of logits for each tool type, using the same methodology as in [64] and then transforms the logits to presence probability scores by a sigmoid operation. Casting the values to binary at a threshold of 0.5 yields the desired result.
3.15.2 Model Training
The provided dataset was split into train, validation, and test sets for the experiment. Training elapsed after 30 epochs using a batch size of 30 frames randomly composed from 3 video clips extracted at 1 fps. Model weight optimization was performed using a Stochastic Gradient Descent (SGD) optimizer (Momentum $\rho=0.9$) with an initial learning rate ($\lambda=7e^{-3}$), decayed by a factor ($\eta=1e^{-4}$) and regularized by an $L_{2}$ norm. Optimal hyper-parameters were selected using the validation set.
The presented SANet model was trained on 4x Quadro RTX 6000 24GiB GPU provided by HPC Unistra Mesocenter managed by the University of Strasbourg, France.
3.16 TeamZero
Members: Muhammad Bilal, Taofeek Akinosho, Adnan Qayyum, Massimo Caputo, Hunaid Vohra, Michael Loizou, Anuoluwapo Ajayi, Ilhem Berrou, and Faatihah Niyi-Odumosu
TeamZero has a multi-disciplinary focus that involves experts from academia as well as from surgical healthcare. In this challenge, TeamZero participated in Category 1, where the task was to classify different surgical tools appearing in the field of view of endoscopic surgical videos. The objective of the team was to develop a systematic methodology to solve multi-label (multi-class) classification problems by harnessing the capabilities of ensemble models. The methodological approach that was followed by TeamZero for their entry in the SurgToolLoc challenge along with the discussion of findings from their experimentation is presented below.
3.16.1 Method Description
Methodology proposed by TeamZero for robust detection of surgical tools consists of three key components: (1) intelligent data preprocessing; (2) model training; and (3) model finetuning (as depicted in Figure 44). Details about the various processes of each component are presented below.
Figure 26: TeamZero: Proposed methodology for robust detection of surgical tools with noisy labels.
Data Description and Pre-processing
TeamZero performed intuitive data preprocessing to remove different types of noises present in the data (as shown in Fig. 27 (a)). This ensures good quality data and ultimately helps develop efficient models. The following are the key steps involved in preprocessing the surgical videos.
  1. 1.
    Frames Extraction: Firstly, they extracted frames from surgical videos. Each video clip contains an average of 1,800 frames, resulting in a collection of over 4 million images with dimensions of 1280x780 pixels. The sheer volume of data, exceeding 1TB, presents a challenge when attempting to conduct meaningful experimentation using the entire dataset. To address this issue, the team initially extracted 10 random frames from each video for preliminary experimentation and hyperparameter tuning purposes. Subsequently, they extracted frames at a rate of 1fps (30 frames per video) for training the final models.
  2. 2.
    Dynamic Cropping using Segmentation: TeamZero performed dynamic cropping of frames to extract the region of interest (ROI) in the camera field of view (FoV). Specifically, they trained a segmentation network (UNet model) using a small dataset that was randomly selected from the data. They used the Prodigy tool and a custom algorithm for the annotation of randomly selected frames from half of the surgical videos. They selected one frame per video from the annotated dataset for creating data for the training segmentation model that was trained for binary image segmentation into foreground and background. Their learned segmentation network efficiently removes (1) the user interface control panel that appears at the bottom; (2) the disclaimer notice at the top; and (3) the black borders appearing on the left and right sides of the videos, as these extra pixels can only contribute towards prolonged network training and wasted computation. A visual example of their dynamic segmentation-based cropping is shown in Figure 27 (b).
  3. 3.
    Black Frames Isolation: In the dataset, a few videos contain entirely black frames without any visual information, while they have also received labels due to extrapolating their parent ones. They have isolated such frames for fixing their labels.
  4. 4.
    Labels Standardisation: TeamZero used lambda expressions to clean label strings, i.e., removing different symbols that include brackets ([]), quotes (“), hyphens (-), and slashes (/). Also, the underscore (_) symbol was replaced with white space.
  5. 5.
    Intelligent Splitting: TeamZero realised that randomly splitting frames into the training and validation set will allow frames of one video to end up in both datasets which will also lead to data leakage. Therefore, they devised a logic to intelligently split the dataset into training and validation sets, where the frames from videos are allocated to either training or validation sets unless the labelled tool combination has only one video in the dataset. This ensures that each tool combination is present in both training and validation sets, as there are eight tool combinations with just one video.
  6. 6.
    Label Extrapolation: The dataset provides one label for each video, while the expected algorithm is required to predict labels for instruments’ presence for all the frames in the video. Therefore, video labels are required to be extrapolated for frames. This extrapolation introduced a lot of noise in the training labels. There are many examples where the label might indicate the presence of three instruments but only two or fewer tools can be seen in the given frame. This is often the case when the surgeon moved the tool out of view even though it is installed on the robotic system. Learning from noisy labels is one of the unique challenges this competition presented.
  7. 7.
    Label Fixing: There are many cases in which video frames are entirely black. This happened when the surgeon uninstalled the camera or wanted to move it to another arm. These are totally or partially black videos. Instead of assigning actual labels to such frames, TeamZero decided to introduce a new ’blank’ category and assigned it to all four arms for black images ([blank, blank, blank, blank]). They used the PIL library’s getextrema method to identify such void images.
Figure 27: TeamZero: Examples of noisy labels (a) and cropped Images using Segmentation Learner (b).
Model Training Strategy
As illustrated in Fig. 44, nine different steps are involved in the development of the model(s) for the detection of tools. Here TeamZero provides the description and intuition of using each method. Their modelling approach involved experimentation on data augmentation, model stability checking, customising loss function, grid search optimisation for architectural search to design ensemble learners, progressive resizing, label smoothing [65], fine-tuning with weighted data loaders, and fine-tuning with the cleaned dataset.
The team carried out exhaustive experimentation to understand three learning strategies for modelling similar multi-classification problems from data with noisy labels. The meta-strategies under consideration included binary classification, multi-label classification, and multi-target classification. The team started with the most intuitive approach of binary classifiers to detect the tool’s presence or absence from surgical images. Binary classifiers train models that can learn to perform one task with great precision such as the presence or absence of bipolar forceps. An obvious deployment challenge in this meta-strategy phase is the huge computational time. The system needs to run and collate predictions from several models (in this case 14 models; one for each tool) to make the final prediction. Most importantly, this meta-strategy cannot learn unique intrinsic relationships among objects which is key in tools usage patterns in the surgery. For example, certain tools are often used in tandem with each other. Since binary classifiers learn and predict in isolation, the resulting model fails to learn similar intrinsic usage pattern features. They then tried multi-label classification in which a model learns to predict multiple labels at once. This strategy compromises the learning abilities of binary classifiers in lieu of learning intrinsic tool usage features, but it is susceptible to learning noisy and spurious correlations. To deal with the limitation of both meta-learning strategies, they then employed multi-target classification which is predominantly used to fuse diverse data types in modern deep learning (DL) applications such as images and text. This strategy allowed the learner to learn some interesting features informing robotic arm-specific tools, which are used often or seldom or never with each other.
TeamZero started the experimentation with minimalistic data to fast-track iterations and see what will work for training the multi-class learner. An exhaustive hyperparameter tuning was performed to choose the appropriate data augmentation types, learning rates, and architecture types. Major algorithm families including ResNet, ConvNext, Swin V2, EfficientNets, ResNexts, REGNETX, and ViT were explored. They selected the final model to be an ensemble of pre-trained models from various families through bagging to predict the presence of tools in surgical videos. To allow the model to learn some of the difficult classes, they also employed weighted data loaders and curriculum learning for extrapolating certain scarce classes. They finetuned the ensemble models for several epochs using a weighted data loader to improve learner performance. All experimentations were conducted using the FastAI library in Python language [66].
3.16.2 Preliminary Performance
Results Description
The following are the key observations from their experimentations for surgical tool detection.
Effect of Dynamic Cropping: Their binary segmentation strategy to dynamically crop images for the borders and UI control which were leaking the tool names phenomenally. It increased the model accuracy and generalisation ability significantly as images without leaky labels forced the models to learn the actual representation of instruments rather than some spurious correlations.
Data Augmentations: TeamZero tested three data augmentation approaches including Squish, Crop and Pad. They found that squishing images works better than cropping and padding. In addition, they also checked for the rectangular vs squared resizing. Their baseline algorithm (ConvNext) performed better when used on a dataset which is scaled with a similar aspect ratio to the original dataset which has a rectangular dimension.
Multi Target Classification: TeamZero found that multi-target classification supersedes the simple multi-label classification on this problem. In multitarget classification, the ML model tends to learn to model each robotic arm. This allows the ML model to learn the intrinsic relationship between tools usage patterns. The multi-target version exceeded the detection accuracy over a single multi-label classifier by 5%.
Model Training using Progressive Resizing: TeamZero employed progressive resizing for the training baseline model. The training starts with small images (4x smaller in this case) and completes the training using large images (mostly original image size if that is not too large). They retrained the aspect ratio while resizing as discussed above. The final model reduces the images to $3/4$ of the original input image by keeping the original aspect ratio. They performed 12 epochs for the smaller image experimentation and then used 6 epochs each for two successive resized datasets. The model was able to improve the accuracy by 1.35%.
Using Test Time Augmentations: To improve the generalisability of their model, TeamZero employed test time augmentation (TTA) which reduced the overall error rate slightly in the 3rd digit place after the decimal. In TTA, they created various augmentations of the original image and used them to make a prediction in addition to the original image and then averaged the probabilities. A key reason for TTA to not work in this case is that they squished the input images rather than using a random resize crop where TTA has shown huge improvements. It did not affect the training time but significantly increased the validation and inference time as the model has to go over images n times, which is the number of augmented versions they specified for the learn.tta() method.
Hyperparamter Selection: They used a higher learning rate for first-time training of the models with 12 epochs and then lowered the learning rate in all fine-tuning afterwards. The learning rate of 1e-2 seemed to have worked very well for all the architectures. They used FastAI’s built-in function to inform this decision. A typical learning rate chart can be seen in Figure 29:
Effect of Ensemble Model: The approach that dramatically improved TeamZero’s rank on the preliminary testing leaderboard was the strategy they used for ensembling. Rather than ensembling using typical approaches such as cross-validation, they trained models across different families of architectures. They used the timm library for downloading these models which have been nicely integrated into the FastAI library. They performed exhaustive experimentation to see which model families can learn the given classification tasks with greater accuracy. They found that ConvNext and SWIN outperform all other architectures. Besides, ViTs and REGNETX were also equally performing well. So, they chose to use four models of different families across these four families. They used smaller architectures to choose the right families and then tested their larger versions for ensemble learning. They found that larger architectures were not improving the overall performance but instead were massively overfitting on the given task. So, they decided to use the smaller architectures instead. Tyhe also doubled the weightage of the predictions from the top-performing models. These included Convexts and SWINs.
Figure 28: TeamZero: Learning rate chart showing 5e-3 where ViT model can learn the most on the given dataset.
Error Analysis: Once TeamZero had finalised their ensemble model, they then performed error analysis. They held out at least one video across different label combination types to understand the learning capability of the ensemble model. The first insight they got was that model is learning the wrong labels with high confidence and accuracy as the right ones. The arm where they observed the most noise was USM4 and USM1. The instrument which is predicted largely wrong by the model is Cadiere forceps. This led us to perform label cleaning at the video level. There were many videos where just two or fewer tools were present, but the label said otherwise. They fixed those videos and fine-tuned their models on newer less noisy data which seemed to improve their performance slightly. But the difference was not dramatic as in most of the videos labels were very wrong and this small fix would still not guide the algorithm to learn that difference.
Manual Label Fixing: They then chose 5 videos per label combination to fix labels at the image level. It was an exhaustive exercise. They used Prodigy for label fixing. They created a custom recipe for this purpose to perform this correction. They fine-tuned their model on this dataset but there was no considerable increase in the accuracy of models even after performing label fixing.
Conclusions and Future Work
Classification tends to model the analytical problem using the one-hot encoding of target labels. The model is forced to learn classes with high confidence since they are eventually rounded to 0 or 1. They felt that this is very harmful as the provided data has various types of noise. Many images are mislabelled. In real life, no data is noise-free. Even human-curated labels end up with wrong labels. Label smoothing allows models to relax label boundaries slightly. So instead of 1s and 0s, the loss function is engineered to use a number less than 1 for 1s and a number a bit more than 0 for all 0s in the encoding vector. Retraining this should have worked. They tried fine-tuning their model using label smoothing but they didn’t have enough time due to wasting a lot of time figuring out docker submission issues. They trained their models for 10 epochs but there was no significant improvement, so they discarded the idea. For label smoothing, models shall be trained for longer to pick up confusing cases. In their future work, they plan to use different noise modelling approaches for fixing noisy labels [67] such as iterative label clearing during model training, automatically identifying, and fixing incorrect labels, using Bayesian models, etc. A major limitation of their proposed methodology is that they did not start off with those prevalent strategies.
Figure 29: TeamZero: Learning rate chart showing 5e-3 where ViT model can learn the most on the given dataset.
3.17 Results
The team evaluations for the challenge were performed on a private, hidden test set containing individual files of surgical videos at 1 FPS. The results were generated through the Grand Challenge automated algorithm submission and evaluation system. Below we present a summary of the challenge results.
3.17.1 Category 1: Surgical tool presence detection
For Category 1, the performance of surgical tool presence detection was evaluated with the mean F1-score [68] across all tools. Overall, the teams performed well in this category with some of them obtaining F1-scores above 0.7 (Table 13). It is important to notice that the presented F1-scores take into consideration data imbalance: the scores are weighted based on how often each tool is present in the test data. This was an important consideration because some tools did appear much more often than others in this data set.
Table 13: Results for Category 1
Team Avg. F1-score
HRI_MV 0.7485
HKMV 0.72
NVIDIA 0.7055
ANL-Surg 0.6691
HVRL 0.6605
SK 0.643
TeamZERO 0.6299
VANDY-VISE 0.6206
UKE 0.6203
ITeM 0.5978
CAMMA 0.5744
MM 0.5637
BioMedIA 0.5521
3.17.2 Category 2: Surgical tool localization and classification
Table 14: Results for Category 2
Team mAP
HRI_MV 0.4077
HKMV 0.3726
NVIDIA 0.3058
ANL-Surg 0.0961
SK 0.0247
VANDY-VISE 0.0003
HVRL 0
ITeM 0
For this category of the performance of surgical tool localization and classification was evaluated with the standard COCO dataset evaluation metric [69], i.e. the mean average precision (mAP) of intersection over union (IoU), discretized with values $0.5:0.05:0.95$ at 1 FPS. This challenge proved to be very demanding for most teams. Only the top 3 teams were able to achieve a reasonable performance of above 0.3 mAP on the test set. The remainder of the teams struggled with detecting bounding boxes accurately and could only achieve mAP values between 0 and 0.1 (Table 14).
Unsurprisingly, teams that performed well in Category 2 also performed well in Category 1. Interestingly, lower performance in Category 2 could still translate to reasonable performance in Category 1. This indicates that detection models may be learning image features that do not help with localization. Critically, the method that distinguishes higher from lower-performing results in Category 2 is pre-training on some data sets with ground truth tool segmentation labels. This may suggest that “nudging” the model weights to some portions of the weight-space that are informed by tool segmentation, helping learn features that are more relevant to localization.
3.18 Discussion
Alternatives to fully supervised learning, such as self-supervision and weak supervision, continue to advance in the areas of object classification, detection, and localization. This is particularly impactful in surgical data science since large volumes of data can be exploited if noisy or weak labels can be utilized. Our SurgToolLoc 2022 EndoVis challenge provided the community with just such a dataset. With all the trappings of a large, real-world dataset (e.g. variability, imbalanced classes, noisy and weak labels), teams could explore how different model architectures and training strategies could solve a relevant surgical task.
In total, there were 13 complete submissions in Category 1 and nine complete submissions in Category 2. Team reports were included to provide a comprehensive view of the approaches taken. Teams performed relatively well in Category 1 but struggled in Category 2. Good performance in Category 1 generally translated to relatively good performance in Category 2. This was perhaps not surprising since discovering an accurate and robust representation of the tool presence, at least conceptually, seems like a prerequisite for localizing these same tools.
For both categories 1 and 2, supplementing training data with publicly available surgical data (e.g. Cholec80), or building off of previously trained backbones (e.g. resnet) proved to be key in getting high-performing models. Specifically, five out of the 13 teams used a clinical dataset to pre-train the model. For Category 1, three out of the top five teams used the EndoVis instrument segmentation dataset to pre-train the model, including HRI-MV, HKMV, and ANL-Surg. Similarly for category 2, two out of the top three teams used the same dataset, including HRI-MV and HKMV. For Category 2, the use of class activation maps was crucial but only proved moderately successful. On the whole, the challenge was difficult and the proposed tasks remain largely unsolved.
Performance in Category 2 was clearly worse than that of Category 1, but also poor by conventional standards. This was not surprising as only noisy, weak labels were available for training. All other things being equal, we assume supervised training would improve performance on this task. By how much is not clear. It could be that performance of these weakly supervised models would be viewed more favorably in light of the limits of similar models trained with strong labels. This would have important implications given the difficulties and resources required in obtaining strong labels. We plan to pursue this question in future work.
There are a host of topics that can be explored with the dataset from this competition. For example, surgical gesture and activity recognition, and tool-tissue interaction are areas of study that this data set can speak to, albeit with various annotation efforts. Similarly, automated, objective skill assessment is an important area that could potentially benefit from this and similar data sets. We hope that this data set can contribute to many future advances.
In this challenge we encouraged the surgical data science community to tackle the problem of automatically detecting and tracking surgical instruments in endoscopic videos. This had to be accomplished by leveraging only tool presence data, as this data is readily available in robotic-assisted surgery. For category 1, designed to classify tools present in each surgical frame, the participating teams developed algorithms that performed well. For category 2, classifying and detecting the spatial locations of surgical instruments, the algorithms performed relatively poorly. The models used by the teams represent many state-of-the-art classifier algorithms. While these models perform well on natural and conventional image data sets, they failed to exhibit similar performance on these surgical videos. We can only speculate as to the reasons for these shortcomings. For instance, it could be due to the noisy labels; it could be that these surgical images represent an entirely different distribution of data that is more difficult to categorize; it could be that this problem has exposed the limitations of these algorithms; or, it could be that these algorithms simply require more training data. While the lack of sufficient training data is the usual culprit for poor performance, it is worth noting once more that we believe this is the largest publicly available surgical data set to date. Regardless of where the models fell short, it is clear that this problem has not been adequately solved.
3.19 Acknowledgements
Work done by the CAMMA team was supported by French state funds managed within the Plan Investissements d’Avenir by the ANR under references: National AI Chair AI4ORSafety [ANR-20-CHIA-0029-01], Labex CAMI [ANR-11-LABX-0004], DeepSurg [ANR-16-CE33-0009], IHU Strasbourg [ANR-10-IAHU-02] and by BPI France under references: project CONDOR, project 5G-OR.
Work done by the ITeM team was supported by the German Federal Ministry of Research and Education (BMBF) under grants CoHMed/DigiMedOP grant no. 13FH5I05IA, CoHMed/PersonaMed-A grant no. 13FH5I06IA, and CoHMed/PersonaMed-B grant no. 13FH5I09IA.
4 Results and methods from the MICCAI 2023 SurgToolLoc challenge
4.1 SurgToolLoc 2023 Challenge Description
4.1.1 Overview
This subchallenge was arranged as part of the Endoscopic Vision Challenge at MICCAI 2023 and was a continuation of the SurgToolLoc 2022 challenge. However, this time there was only one category of surgical tool detection (the same as category 2 of 2022 challenge). Please see the description provided in Section $3.1$ for more details.
4.1.2 Training Data
The training data used for this challenge was the same as 2022 SurgToolLoc challenge - please see Section $3.1$ for specific details on the dataset.
4.1.3 Testing Data
The format of the testing dataset was exactly the same as that of 2022 Challenge, however, a new set of 108 video clips collected from surgical training exercises were used as the test set. This meant that the algorithm performance numbers (mAP values) for the two years cannot be directly compared. However, we did evaluate the winning algorithms from 2022 in the 2023 test set to be able to compare and see if the algorithm performances improved over the year.
The testing data was a set variable mean (SD) clips in seconds: 687.45 (424.68), downsampled to 1 FPS for inference purposes. The test data was annotated with bounding boxes around the instruments, using a crowd of experienced annotators. The process of creating the 2022 testing set labels was reproduced here; see section $3.1$ for more details.
4.1.4 Submission Process
The submission process for this challenge was the same as 2022 challenge; please review Section $3.1$ for more details.
4.2 Team Submissions
For this year’s challenge, there were a total of 18 teams that had shown interest in participating and downloading the dataset. In the end, only 7 teams submitted full submissions in the single category available. Table 15 shows the participating teams, while Table 16 summarizes the methodologies employed by each team. Team methodological details follow below. Note that these sub-sections were written by the participating teams.
Table 15: Team affiliations and challenge categories - 2023
Team # Team name Institution Country Report
1 AIT King’s College London UK Y
2 ANL-Surg Argonne Nat. Lab. USA Y
3 CAIR-HK City U. of Hong Kong Hong Kong Y
4 HVRL Keio U. Japan Y
5 Jmees Jmees Inc. Japan N
6 LabRen-CUHK Chinese U. of Hong Kong Hong Kong N
7 LozaGera U. of Leeds UK N
8 MapleLab Vanderbilt U. USA Y
9 PUMCH Peking Union Med. Coll. Hosp. China Y
10 SDS-HD German Cancer Res. Ctr. (DKFZ) Germany Y
11 seventeen Beijing U. of Posts & Telecom. China N
12 TUE-VCA Eindhoven U. of Tech. Netherlands Y
13 ZJURealdoctor Zhejiang U. China Y
Table 16: Summary of methodologies - 2023
Team Name Architecture Backbone Data Preprocessing Pretrain Image Augmentation Use Additional Data Loss Output AIT Faster-RCNN + U-Net ResNet OCR for UI detection Y EndoVis2017 Binary cross-entropy Tool bounding box ANL-Surg Detectron2 Faster RCNN + YOLOv8 ResNet Manually annotated frames Team Nvidia’s dataset from 2022 challenge Tool bounding box CAIR-HK WS-YOLO CSPDarknet53 Pseudo-labels Y Tool bounding box MapleLab TernausNet + ResNet-101 ResNet Attention masks for tools + pseudo labels Y Binary cross-entropy Tool bounding box PUMCH OsTrack Transformer tracking model Auto-generated pseudo-labels Custom (see report) Tool bounding box SDS YOLOv5 + SwinTransformer CSPDarknet53 Y Semi-automatically generated bounding boxes Binary Cross-Entropy Tool bounding box TUE-VCA YOLOv8 CSPDarknet53 Filtering Y Pseudo-labeling Symmetrical cross-entropy Tool bounding box
4.3 AIT
Aiming to keep within the spirit of weakly-supervised object detection, our approach tried to make use of only a small amount of human annotated data. We also note that the presence labels provided are often incorrect and the UI is a better source of information as well as providing additional information such as activation and rough positioning which we make use of.
We start with EndoVis2017 tool parts segmentation dataset to train models for binary tool segmentation and binary tool tip detection. We then use this in conjunction with the data parsed from the UI of the SurgToolLoc data to generate a pseudo labeled dataset. We then train our final model on this data, using any SurgToolLoc data which could not be assigned pseudo labels as weak supervision.
4.3.1 Method Description
The EndoVis2017 dataset was used to train models for both binary tool-tip object detection and binary tool segmentation. For the detection model we used a basic FasterRCNN [70]. The detection labels were generated using the tool parts classification maps. The bounding box was drawn around the clevis, if no clevis is visible, it is instead drawn around the tip. As tool part segmentations do not exist for the testing samples, we make our own training and validation splits from the training samples with 6 and 2 clips respectively. For the segmentation model we train a standard UNet [71] on the segmentations provided with the training/validation split following the training/testing splits provided in the original dataset.
The UI present in many of the clips provides richer and more accurate information than the ground truth tool presence labels. We implement a parser using EasyOCR [^12] https://github.com/JaidedAI/EasyOCR for optical character recognition (OCR) to read the tool names and note the order these tools appear from left to right. The ordering of the UI elements representing the tools and the camera can be used to determine if the robot is configured with two tools on the left and one on the right, or visa-versa. We also use a simple color model to determine which tools are currently active, as these are displayed with a pale-blue UI instead of gray.
To generate our pseudo labeled dataset we combine inferences from our two trained models with the information parsed from the UI. We sample the provided SurgToolLoc video clips every 30 frames, to reduce residency in the data, and run our UI parse on these frames If the OCR finds no tool names the clip is assumed to have no UI and is discarded from the pipeline. We then run our binary segmentation and detection models on the frames. The segmentations are processed by keeping connected components above 400 pixels in area. Any frames that do not have exactly three connected components remaining are discarded. The detected bounding boxes are then assigned to one of the three connected components of the binary segmentation based on intersection over union (IoU). We then keep the most confident detection for each connected component. The tools are determined to be either on the left or the right based on where the binary mask meets the edge of the image. We also run RAFT [teed2020raft] optical flow to both the previous and next frame, averaging the forwards and backwards magnitudes of the flow vectors. We average the flow magnitudes over each connected component of the segmentation map to arrive on an average motion for each detected tool, with tools with average motion above a threshold are determined as active. We now use our detect labels (left, right, active) for each bounding box and compare these to the labels parsed form the UI. The correct class label is assigned to the bounding box, if there is a discrepancy, the sample is discarded. This whole process has a success rate of around 6The class distributions for the data can be seen in Figure 30.
To incorporate the discarded data, we use the classification logits of the detections to produce a multi-label prediction for the image. Specifically, the classification logits of the detections are summed together, weighted by the confidences. A loss can then be calculated between the prediction and the presence label provided in the original SurgToolLoc data.
Figure 30: Class wise distribution for the original SurgToolLoc labels and the generated pseudo labels. The pseudo labels are also presented re-scaled to better compare the distributions with the SurgToolLoc labels
4.3.2 Model Training
We again use a FasterRCNN, initialized with the weights from the model pretrained on EndoVis2017. We batch together images from the weak and the strongly labeled datasets, employing class balanced sampling to ensure classes are more evenly represented. We apply random photo-metric and spatial augmentations to all the images. The strong supervision loss is calculated as the standard FasterRCNN loss summed and the weak supervision loss calculated using binary cross entropy. The model parameters were optimised using stochastic gradient descent with learning rate of $1$$\times$$10^{-3}$ , moment $0.9$ and weight decay of $1$$\times$$10^{-4}$. We reserved 10% of our pseudo labeled data to use as a validation set for early stopping.
4.3.3 Preliminary Performance
Training on our pseudo labeled data with additional weak supervision we were able to consistently achieve around 0.33 mAP on our hold-out validation data. However, submission to the testing server yielded a much lower score of 0.11 mAP suggesting our model was badly over-fitting to some aspect(s) of the pseudo labeled data. An ablation study showed our weak supervision, provided only a modest improvement of just 0.03 mAP on our validation data.
While our pseudo labels generation provided accurate detection labels, too much of the data was discarded to ensure correctness of the retained data, leading to a small overall dataset. Furthermore, it is likely that the classes which were not present during the training of our binary object detection model were underrepresented in the final dataset. Ideally, weak supervision would provide a means to overcome these deficiencies, however we found our method was insufficient to effectively incorporate the information in the weakly labeled data. Perhaps, given more time, a more effective weak supervision approach would have been settled on.
4.4 Argonne National Laboratory - Team ANL-Surg
Team ANL-Surg includes Assistant Computational Scientist Neil Getty from Argonne National Laboratory (ANL), and Ruoxi Zhao, a recent graduate in Applied Mathematics from UC Merced. The team was primarily interested in Category 2. Motivated by the potential of weakly-supervised learning methods, ANL-Surg trained Detectron2 Faster RCNN [70] and YOLOv8 [72] on custom dataset for localizing, detecting and tracking each surgical instrument.
4.4.1 Method Description
Team ANL-Surg’s goal was to train the Detectron2 Faster RCNN model on approximately 2500 manually labeled frames, and utilized the Mixformer [73] model for tracking the clevis of each tool starting from the manually selected frame in the clips, then performed backward and forward tracking at 30 frames per second (FPS). The team also trained the YOLOv8 huge model on the extended dataset, with a mean Average Precision (mAP) of 59.7% on their validation holdout. The motivation of choosing the described models was to get accurate results of tracking and recognizing the surgical tools in the frames, without excessive human intervention and manually labeling all the frames in the training dataset.
Their approach began with the careful annotation of data. They manually annotated approximately 1500 frames from the competition dataset. These frames were selected halfway through the video clips, with a focus on the underrepresented surgical tools in the dataset, such as clip applier, permanent cautery hook/spatula, stapler, suction irrigator, tip-up fenestrated grasper, and vessel sealer. The reason for building this dataset was to increase the visibility of the underrepresented tools to the model, which the model tended to perform poorly in the previous training sessions due to the larger size of such tools. In addition, they edited team Nvidia’s dataset from MICCAI 2022 SurgToolLoc challenge that had around 1000 labeled frames, to be more accurate in capturing the clevis of the surgical tools, adapting to the competition guidelines for this year’s challenge. The foundation of their methodology lay in model training. The team initiated their efforts by training a Detectron2 Faster RCNN model using those 2500 manually annotated frames. This model served as their baseline for subsequent improvements.
4.4.2 Model Training
To enhance the temporal understanding of tool localization, the team employed the Mixformer object tracking model. This tracking model was used to follow labeled clevis of the tools in both forward and backward directions within the video clips, operating at a frame rate of 30 FPS. Tracking was initiated from the manually annotated frames. After using the trained model to infer tool detections of all the 30 FPS frames, the team merged tracking boxes generated by the object tracking model with prediction boxes produced by the Detectron2 model. This integration allowed them to rename predictions based on tracking data, effectively propagating manual labels throughout the entirety of the video clips. To reduce redundancy in the dataset, image hashing techniques were implemented. This step aimed to identify the frames that exhibited high similarity, remove frames where tracking differed from prediction by threshold, mitigating duplication and streamlining the dataset for training and evaluation.
Figure 31: Label frequency of manually annotated tools.
Figure 32: Example image showing tracking propagated labels (red) and overlapped predicted class.
The model refinement process continued with the training of a YOLOv8 model. This model was trained on the expanded dataset, including a validation holdout. Its primary focus was on detecting frames containing three surgical tools, which the frames would have more potential of matching the three global labels in model predictions. For enhanced reliability and accuracy, the team enforced strict label matching criteria. Predicted labels were required to match global labels with high model confidence. Frames that did not meet this criterion were removed from consideration. They implemented a confidence threshold of 70% and max detection to 4 to classify detections as trustworthy, acknowledging the model’s certainty in its predictions.
For data augmentation, random 30% rotation and 10% shearing was used. The model was trained up to 300 epochs with a learning rate of 5e-4 and a batch size of 102 images. 6 V100 GPUs with 32gb memory were utilized for training.
4.4.3 Preliminary Performance
The comprehensive methodology yielded promising results. Initially, the YOLOv8 model trained solely on manual labels achieved a validation mAP of 42.7% on the holdout set. As they progressed, the model trained on tracked images, consisting of 136,237 frames, demonstrated an improvement, achieving a validation mAP of 49.1%. Notably, further advancements were observed as the model was trained on tracked and predicted frames, specifically those containing three tools. This iteration achieved a validation mAP of 55.7%. However, the most significant leap in performance occurred when the team trained the model on the entire dataset, incorporating both high and low threshold detections, resulting in the highest validation mAP of 59.7%. Finally, the ultimate model, trained on the complete dataset without a holdout, comprised 432,918 images and exhibited substantial enhancements in surgical tool localization accuracy. These results highlight the effectiveness of the methodology in progressively improving model performance.
Figure 33: Confusion matrix for all classes.
In conclusion, after testing out different methods and models, the team achieved a validation mAP of 59.7% with the YOLOv8 model. The use of the Mixformer object tracking model helped the team to gain more training data by tracking the tools from one labeled frame, significantly reducing the time and labor of manual labeling medical dataset. Additionally, the use of confidence thresholding helps identify untrustworthy detections, enhancing the reliability of the localization model. In future research, the team is interested in focusing on the frames where one or two tools were present, which usually contains larger tools such as clip applier, stapler, and tip-up fenestrated grasper. The team is also interested in the unsupervised methods for this challenge, such as training DINOv2 [74] on the training data to learn the features of the surgical tools in a self-supervised way.
4.5 MapleLab
Team Members: John Han, Ayberk Acar, Jumanh Atoum, Yinhong Qin, Jie Ying Wu
The general intuition of team MapleLab’s model was that by seeing many examples of different tools and their combinations, a model would be able to associate which tools are which while learning the image labels during training. Therefore, the model could correctly classify tools by recognizing the overlap of tools across different video samples and distinguishing each tool from the rest.
4.5.1 Method Description
The MapleLab team developed a pipeline that did not require any additional manual labeling and learned the tool classification and localization from the complete surgical scene. The inputs of the model were the color image, segmentation mask generated from TernausNet, and their element-wise multiplication (referred as the attention mask).
The application consisted of three stages: data preparation and bounding box initialization, training, and reiterated inference for classification.
In the first stage, the team employed TernausNet [26], a binary segmentation model trained on the EndoVis 2017 dataset [25], to generate tool masks from the input RGB image. These two images were multiplied and Gaussian blurred to create an ”attention mask.” These three images (the original color image, the binary segmentation mask, and the attention mask) were cascaded to create a 7-channel tensor, which is used for the next two stages. In addition, the method obtained bounding boxes for each tool by identifying the rectangle around each connected component of the segmentation mask. Because the evaluation metric requires bounding boxes around the clevis of each tool, each bounding box was scaled down with a manually-selected scale factor of 0.6.
The goal of the training stage was to produce a network that could correctly identify which tools were presented in a frame, using only the tool presence labels. Their model, a fine-tuned ResNet-101 model [75] pretrained on the ImageNet-1K dataset, took the input of the 7-channel tensor created in the first stage and predicted the tools that existed in that frame in one-hot encoding format. Figure 43 shows an example of the input 7-channel tensor and model behavior.
Figure 34: A fine-tuned ResNet-101 model takes the concatenated 7-channel image as input and outputs the tool presence vector.
In the final stage, every bounding box was labeled by reiterated inferences with the trained ResNet model. A score vector $y$, which represented the model’s confidence values of what tools exist in the frame, was first extracted. Next, all tools in the binary segmentation mask and attention mask were masked out except for the targeted tool to be classified. The model received this new set of images and generated another score vector $y^{\prime}$. These two score vectors, $y$ and $y$’, represented the model’s confidence values of what tools were present in the original set of images and the altered set of images. The classification label of a tool was determined to be the maximum of the values of $y^{\prime}$ where there were positive confidence values from $y$. In other words, the most confident label was selected after the other tools had been masked out from the input data. This procedure was repeated for all tools to assign class labels to each bounding box. Figure 35 displays the workflow of the classification process, and Figure 36 shows some example output images.
Figure 35: In the final reiterated inference stage, predictions are made on each bounding box via masking out the other tools.
Figure 36: Qualitative Results
4.5.2 Model Training
To improve the quality of training data, a morphological opening and closing was applied to the segmentation mask to remove noise and close holes after running TernausNet. The images with poor segmentation masks were removed, via comparing the size of the segmentation blob with certain thresholds. Segmentation blobs with areas $>0.02\times\textrm{total image area}$ were kept and segmentation masks with more than 40% predicted foreground pixels in the image were removed. Furthermore, the team also down-sampled given images to the size of $512\times 512$ with 0 mean and unit standard deviation. The train-test split was 70-30. The initial learning rate was set to 1 changing at every epoch with cosine annealing and the optimizer was stochastic gradient descent (SGD). The loss function was binary cross entropy. An NVIDIA RTX4090 GPU was used for the model training.
4.5.3 Preliminary Performance
During the training stage, team MapleLab’s model had an accuracy of 98% in correctly identifying the present tools. The model got a final test score of 0.0007 mAP (mean Average Precision). The low accuracy of the model may be due to the model’s inability to correctly identify which tools are which from the team’s approach, as well as the fact that the bounding boxes were not refined specifically to be ROIs of the clevis of each tool.
While the team’s method acquired high accuracy in the initial training phase, results could be improved by the selection of frames with tool presence matching the provided labels, since the data did not always have all the tools in the field of view. Finally, overlapping tools in the camera view created a challenge for our pipeline, since the model relies on accurately separated tools from the binary segmentation.
4.6 PUMCH
  • Team name: PUMCH
  • Members: Surong Hua, Lu Ping, Wenming Wu
  • Research field: Computer Vision
  • Institution: Peking Union Medical College Hospital
  • City: Beijing, China
  • motivation and plan: At first glance, we were thinking of weakly supervised learning. But then we found that the video tags provided for training only contain information on whether a certain instrument exists in this video, it was not certain that whether the instrument appears in every frame, since it might be partially or completely obscured, which reduced the performance. We also noticed that almost all the winning teams last year were based on semi-supervised method. Therefore, we proposed a framework by combining semi-supervised learning, object detection, and object tracking together.
4.6.1 Method Description
Data Processing
See Figure 37 ,in this framework, we adopt the following three methods to process data. Firstly, for frames without labels, we use the MAE[76] method to train the feature extraction ability of the backbone network. Secondly, we referred to last year’s champion method to merge and process consecutive video segments to get sequence information, and then manually annotate the target in multiple stages of the video. Then, we used a transformer-based target tracking model (ostrack)[77] to generate pseudo labels. Thirdly, we use tools presence labels to conduct cross validation and manual correction. Then, the model is continuously looped and trained to update pseudo labels. Ultimately, we obtained 30,000 pseudo labels of data.
Figure 37: Data Processing
Architecture Design
Specifically, As Figure 38 shows, OStrack[77] in our strategy consists of the following four steps: (1) The template and search region undergo patch embedding and spreading to recover the token, which is added with the positional encoding and spliced to form the overall token. (2) The overall token is fed into the encoder, in which the similarity between each token and the template token is calculated, and the foreground token is retained while the background token is rounded off in descending order. (3) Token padding is then performed to complement the rounded tokens with zeros and output to the head. (4) The head section outputs the classification score, the predicted offset value to compensate for down-sampling quantization error, and the normalized bounding box size, respectively.
Figure 38: OsTrack Pipeline
Model Loss Function
To train the one-stage object detector, we adpoted a dynamic soft label assignment strategy based on SimOTA [78]. As RTM-DET [79]does, we used the IoU between the predictions and ground truth boxes as the soft label to train the classification branch, used the logarithm of the IoU as the regression cost instead of GIoU used in the loss function, used a soft center region cost instead of a fixed center prior to stabilize the matching of the dynamic cost.
Post Processing
We adopted Kalman filtering approach based on weighted boxes fusion (WBF) [80] to achieve a higher detection rate.To be specific, the implementation of WBF has the following steps: (1) The prediction frames of both the detection and the tracking model are normalized by coordinates and confidence, and clustered into a single set B and sorted in descending order using the confidence of the detection frames. (2) two empty lists are initialized, list one is used to hold clusters of detection frames for each target and list two is used to hold fusion frames for each target. (3) The algorithm would iterate through the predictions in B and calculate Intersection over Union (IoU) with each fused frame in list two. (3) When IOU is greater than a manually set threshold “T”, the clusters are considered to be in the same class, and coordinates and confidence of the fused frames are calculated. On the other hand, when IOU is smaller than T, the new box is defined as a new cluster center and is added to the tails of list one and list two. Then the iterative traversal for all the boxes in B is repeated. Eventually the box in List two would be the result.
4.6.2 Model Training
Long Tail
As we noticed that our test metric is MAP, and we realized that the 14 low frequency instruments have long tail distributions, which will lead to the model overfitting,see Figure 39 .To avoid this, we utilized the tools presence labels to generate pseudo labels in similar size by instrument categories.
Implementation Details
We divided all the data into a training set and a validation set by 7:3, and then conducted model experiments, data augmentation, category balance, and ensemble experiments, respectively. For every experiment, we trained our models using AdamW optimizer with an initial learning rate of 0.004, momentum 0.9, and a batch size of 64 for 100 epochs on the training set and validated them on the validation set. We resize the input images to 640 x 640 and the learning rate is annealed following the poly learning rate policy. All experiment were trained on 8 NVIDIA A100 GPUs. We evaluate the model performance on object detection and instance segmentation by bbox AP.
Figure 39: Long Tail
4.6.3 Preliminary Performance
Results
Tab1-Tab3 show which modules perform best on the current validation set.Then, we choose RTM-DET [79] as our detection model, cbnet-swim [81] as our backbone, auto-augment as data augmentation, and adopt category balancing strategy. Finally, we adopted the integration of RTM-DET and the fine-tuned tracking model as our final submission.
Model Size Bbox_map
Cascade R-CNN-Detectors (894,682) 63.6
RTMDet-l (864,864) 70.6
YOLOv8-l (864,864) 69.1
Co-DETR-swin-s Multi scale 64.6
Table 17: Model Experiments
Backbone Bbox_map
cspnext 70.6
Convnext 70.6
Convnext-v2 70.9
swintransformer_7 70.6
swintransformer_12 71.1
Swintransformer_MAE 66.0
resnet 63.6
convnext 70
Cbnet_swin 71.5
Cbnet_res2net101 70.1
Data Augmentation Bbox_map
Mosaic_mixup 70.6
Mosaic_mixup9 70.8
Fog-Blur-Noise and channelshiift 71.2
RandomBrightnessContrast MotionBlur ShiftScaleRotate 71.1
Autoaugment 71.2
convnext 70
Cbnet_swin 71.5
Cbnet_res2net101 70.1
Class Balance Bbox_map
w/o 70.6
w 71.4
Table 18: Main Experiments
Model Fusion Bbox_map
rtmdet-siamfcpp wbf_mean 65
rtmdet-OStrack wbf_mean 66.7
rtmdet-OStrack_FT wbf_mean 67.9
rtmdet-OStrack_FT wbf_max 64.2
Table 19: Ensemble experiments
Conclusion and Discussion
We have adopted the method of integrating detection models with tracking models. Although it has shown effectiveness in offline evaluations, it has not shown greater advantages in online evaluations. We believe that the online test set may be more complex, and our tracking model does not perform well. Compared to the previous semi supervised schemes, we have introduced self supervised learning, optimized the detection and tracking models, and considered class imbalance, which are still directions for future optimization.
4.7 Team Name: CAIR-HK
4.7.1 Method Description
Due to varying instrument visibility across frames, relying solely on video-level labels can induce substantial noise. Developing models capable of effectively detecting surgical instruments necessitates designing pseudo-label generation algorithms to maximize the utilization of available labels. Many existing approaches employ manual annotation to supervise models, but the vast number of videos in the dataset renders manual labeling extremely laborious and non-autonomous. Our approach substantially reduces the requisite human annotation effort while achieving an excellent compromise between manually labeled data volume and model performance.
In this work, we proposed a Weakly Supervised Yolo Network (WS-YOLO)[82] for Surgical Tool Localization in Endoscopic Videos. Specifically, we preliminary trained an object detection model to distinguish the 3 parts of surgical tools (shaft, clevis, and tip) using the annotated public dataset (Endovis [83]). Then, we sampled the 24695 video clips provided by the challenge [84] at a frequency of 1/30 and performed inference on the sampled dataset using the trained model, obtaining an initial pseudo-label dataset via the Initial Label Filter strategies. Subsequently, we trained a new model with surgical instrument detection capabilities based on this initial pseudo-label dataset. We then selected new pseudo-labels by screening the inference results from this and previous models, leveraging Multi-round Label Filter strategy. Generating a new pseudo-label was iterated multiple times to elevate the performance. Our approach significantly diminishes the necessary human annotation labor while striking an optimal balance between the quantity of manually annotated data and detection performance.
4.7.2 Method
Figure 40: The overall pipeline of our proposed method.
Fig. 40 presents the overall pipeline of our proposed method. Initially, we transformed the semantic masks in the Endovis [83] into labels for object detection. We trained a model $Det_{parts}$ based on YOLOv8 [85], capable of localizing the shaft, clevis, and tip, as depicted on the left of Fig. 40. We sampled every 30th frame from each video clip for the dataset. We performed inference with the trained model $Det_{parts}$, on the sampled dataset, obtaining corresponding bounding boxes (localizing the shaft, clevis, and tip in each sampled frame).
Figure 41: Pseudo-code of Initial Label Filter Algorithm.
Due to varying instrument visibility across frames, relying solely on video-level labels can induce substantial noise. Developing models capable of effectively detecting surgical instruments necessitates designing pseudo-label generation algorithms to maximize the utilization of available labels. Many existing approaches employ manual annotation to supervise models, but the vast number of videos in the dataset renders manual labeling extremely laborious and non-autonomous. Our approach substantially reduces the requisite human annotation effort while achieving an excellent compromise between manually labeled data volume and model performance.
Figure 42: Pseudo-code of Multi-round Label Filter Algorithm.
Since the bounding boxes output by $Det_{parts}$ did not contain category information of surgical tools, we adopted an Initial Label Filter approach to match the outputs of $Det_{parts}$ with video-level labels, as illustrated in Fig. 41.
Given that each video-level label contains at least three tool categories arranged in the order of the robot’s three arms (from left to right), based on this prior knowledge, we posited that in most cases the tools from left to right in the surgical scene should correspond one-to-one with the video-level label order. Thus, in Algorithm 1, we first sorted the bounding boxes output by $Det_{parts}$ from left to right. In the initial label filter, we only selected frames that concurrently detected three clevises or three tips to generate pseudo-labels. We designated ’monopolar curved scissors’, ’tip-up fenestrated grasper’, ’suction irrigator’, ’stapler’, and ’grasping retractor’ as special classes requiring tip detection, such that frames were only included in the pseudo-label dataset if $Det_{parts}$ concurrently detected three tips when the label belonged to these categories.
After generating the preliminary round pseudo-labeled datasets, we trained a detection model called $Det_{tools}$ using YOLOv8 [85]. Similarly, we performed inference on the sampling dataset using the trained $Det_{tools}$ model to obtain detection boxes $BBox_{tools}$ corresponding to instruments in each video frame (localizing all instruments in every frame). Since both $Det_{parts}$ and $Det_{tools}$ contain positional information of instruments in the video frames, but the first round of pseudo-labels had considerable noise, $Det_{tools}$ had lower accuracy. Therefore, we adopted a multi-round label filtering approach to select high-quality pseudo-labels, as shown in Algorithm 2 (Fig. 42). For each detection box $bbox_{tool}$ from $BBox_{tools}$ that corresponds to a particular instrument category, we checked if there exist any $bbox_{part}$ with category tip that has IoU $>$ 0.8 with $bbox_{tool}$, and any $bbox_{part}$ with category clevis that has IoU $>$ 0.8 with $bbox_{tool}$. Only when all $bbox_{tool}$ in a frame satisfy these conditions, the pseudo-label is considered trustworthy.
4.7.3 Model Training
The training strategy of basic detection model is following the default setting of YOLOv8, and the iteration strategy of training and data pre-processing are illustrated above. We take 1 Tesla A100 with batch size set to 8 to train 1000 epochs.
4.7.4 Preliminary Performance
In the preliminary testing, we obtained a mean average precision (mAP) of 0.043 using the initial round of pseudo-labeled data generated by Algorithm 1. In addition, we manually annotate a few video clips to figure out the problem of data unbalanced ($clip_{5670}$, $clip_{5672}$, $clip_{5675}$, $clip_{5985}$, $clip_{5986}$, $clip_{5987}$, $clip_{16647}$, $clip_{16650}$, $clip_{16652}$, totally 549 images are manually annotated, where $clip_{i}$ denotes the $i_{th}$ video clip in dataset).
Table 20: The performance of WS-YOLO in terms of iterations.
Iteration $round_{0}$ $round_{1}$ $round_{2}$ $round_{3}$ $round_{4}$
mAP 0.043 0.109 0.135 0.147 0.157
In the process of filtering high-quality tags over and over again, our algorithm creates a positive feedback mechanism that makes performance improve with each iteration. As shown in the Table. 20, through iterative refinement of the pseudo-labels using Algorithm 2, we achieved mAP scores of 0.109, 0.135, 0.147, and 0.157 after successive rounds, demonstrating the effectiveness of the multi-round pseudo-label filtering approach for improving model performance.
4.8 DKFZ - Team SDS-HD
SDS-HD team members: Amine Yamlahi, Piotr Kalinowski, Dominik Michael, Tim Rädsch, Marco Hübner and Lena Maier-Hein.
SDS-HD was motivated by developing an innovative semi-automatic annotation pipeline leveraging the recently introduced keypoints tracking algorithm.
4.8.1 Method Description
Our approach focused on the development of an innovative annotation pipeline, leveraging the recently introduced keypoints tracking algorithm TAPIR  [86]. More specifically, we addressed the absence of bounding box annotations in the training data by a pipeline based on reliable keypoints tracking. To this end, we designed a semi-automatic annotation pipeline, comprising the steps of manual keypoint annotation, automatic keypoint tracking,automatic conversion of keypoints into bounding boxes, and filtering of bounding boxes. The semi-automatically generated bounding boxes were used to train a multi-label instrument detection model. Multi-class classification was achieved with a Swin Transformer model  [87] which received the (cropped) bounding boxes, generated by the detection model, as input.
We propose a two-stage pipeline consisting of a YOLOv5 detection model  [88] applied to individual frames and a Swin Transformer classifier, individually applied to the preprocessed bounding boxes of the detector. The core innovation of our approach is related to how the training data was assembled: Training of the detection model: The semi-automatic annotation of the training data with bounding boxes was achieved with TAPIR for keypoint tracking. Based on a manual initialization of keypoints, the track any point (TAPIR)  [86] model was used to transfer the keypoints to the subsequent frames of the same video. Subsequently, the keypoints were converted to bounding boxes, labeled with the corresponding instrument. After filtering of implausible results, the subset of frames with plausible bounding boxes was used to train an ensemble of multi-label YOLOv5 detection models. The trained ensemble was then applied to an additional set of clips to enlarge the final training set for the classification model. Training of the classification model: For training of the classifier, the data with which the detection model was trained was converted to a set of cropped bounding boxes resized to 224x224 pixels with corresponding instrument label. The same workflow (detection, filtering, cropping) was applied to a set of external videos from five sources [25, 19, 89, 90, 91] disregarding any existing labels from these datasets.
Figure 43: Proposed training workflow: (a) Keypoints tracking-based bounding box generation. (b) Classification based on cropped bounding boxes.
4.8.2 Model Training
Data
Semi-automatic annotation of training data: We selected 1400 clips (6 %) of the provided training data for semi-automatic annotation. Criteria for selection were class prevalence as well as diversity in motion, number of instruments and surgical scene diversity. Incorporation of External Data: To enhance the provided training set with five external detection and segmentation datasets  [25, 19, 89, 90, 91] representing three distinct surgical procedures: nephrectomy, prostatectomy, and gastrectomy. These were used to enhance the variety within our training dataset.
Bounding box generation using TAPIR
For 600 out of the 1,400 clips selected for semi-automatic annotation, we implemented the following procedure:
(1) Initialization: We first identified the first frame in which the relevant parts of at least two of the tools were visible. The clevises of the instruments were then annotated with keypoints and instrument label. We generated our annotation protocol based on the shared information of the challenge organizers. Following best practices  [92], we added additional images for the annotation classes and discussed ambiguous scenarios, in order to apply a consistent annotation style regardless of the annotator. Our annotation protocol is presented in the shared code repository.
(2) TAPIR-based tracking: All keypoints set manually were subsequently tracked in the entire clip using TAPIR. We displayed example images for the relevant classes directly in our annotation tooling to enable an easy access for the annotator to the annotation protocol information.
(3) Bounding box conversion: Each set of keypoints corresponding to one instrument was converted to a bounding box.
(4) Filtering: To overcome annotation uncertainties of TAPIR, resulting in keypoints “jumping” from one tool to another, or in moving to the background, we removed frames that met one of the following criteria:
  • Area Rule: At least one of the bounding boxes in the frame is substantially smaller (20%) or bigger (80%) than the corresponding initial, manually annotated box
  • Narrowness Rule: The frame contains at least one box with a height-to-width or width-to-height ratio less than 0.2.
  • Jumping Points: It features bounding boxes based on keypoints that transition from the instrument to the background. This is achieved by implementing a minimum distance margin of 20 pixels between two consecutive frames.
(5) Manual refinement: As a final step of quality control, any noisy annotations that managed to pass through the filters were manually removed. Optionally, the original keypoints could be refined, leading to a repetition of steps (2)-(4). This process applied to the 600 clips yielded 114134 frames, which were used to train a multi-label YOLOv5 detection model. The trained model was applied to further 800 clips from the training data as well as 6500 frames from the five external datasets. After manual cleaning, the process resulted in a total of 268020 frames annotated with bounding boxes and corresponding instruments.
Instrument detection
Instrument detection was accomplished by training three YOLOv5  [88] models: two L6 (with and without external data respectively) and one X6, utilizing annotations generated by the keypoints tracking model. The training process employed an initial learning rate of 0.003 and a final learning rate of 0.004, stochastic gradient descent (SGD) optimizer and Binary Cross-Entropy (BCE) loss function. Augmentations applied during training encompassed rotation, flipping, adjustments to hue, saturation, and value, in addition to mosaic augmentation techniques. An ensemble of the three models, created using weighted box fusion  [80], yielded the final model.
Instrument classification:
The performance of the multi-label YOLOv5 detection model in terms of localization was adequate; however, classification of the instruments occasionally posed challenges, particularly when dealing with instruments that shared similar clevises, such as the cadiere forceps and the prograsp forceps. To address this limitation in our pipeline, we introduced a Swin Transformer  [87] second-stage classifier to improve the classification results generated by the integrated classifier of the detection model (Fig. 1). This classifier operates on the cropped bounding boxes generated by the detection model, with the added step of enlarging the bounding boxes to encompass the instrument’s tip. This enlargement provides additional context and aids the classifier in distinguishing between similar instruments more effectively. The model underwent training for 20 epochs, employing a cosine annealing scheduler that initiated with a learning rate of 2e-4 and concluded at 2e-5. The training process employed a binary cross-entropy loss function and incorporated augmentations such as flips, rotations, and color jitter.
All the models were trained with Nvidia RTX 3090 GPUs.
4.9 Team Name: TUE-VCA
Team member: T.J.M. Jaspers
4.9.1 Method Description
Figure 44: Team TUE-VCA Schematic representation of the proposed workflow.
Figure 44 presents a schematic representation of our proposed methodology used for the SurgLoc2023 challenge. Our method comprises several steps that culminate in the development of our final detection model. Initially, we work on preprocessing the data at a classification level, minimizing the amount of label noise present in the data. Following this, we transition from classification to detection. The refined detection labels serve as the final data for training our YoloV8 detection model.
Classification preprocessing
1) To initiate the process, we adopted the data-processing pipeline recommended by Team NVIDIA in the SurgLoc 2022 challenge, which employs a strategy of dividing the video chunks based on the differences in their hash values [84]. This preliminary step of preprocessing facilitated the efficient organization of the video data. We decided to construct a small validation set, deliberately including at least one instance of each instrument type. The remainder of the video data was allocated for training purposes. Given the constraints of computational resources, we opted to extract 2 frames from each video chunk. This resulted in a total of 47,227 frames for training and 1,073 frames for validation.
2) Initially, our dataset encountered a challenge due to the occasional invisibility of the tool’s clevis within the frame. This resulted in a discrepancy between the automatically generated labels and the tools visible in the frames. To address this, we employ a segmentation model, leveraging the dataset of the EndoVis17 challenge, which is specifically tailored for robotic instrument segmentation [25]. The segmentation model adopts the Unet architecture with a Mit-b4 encoder [71, 93], and is trained using the cross-entropy loss function with the Adam optimizer, employing a learning rate of 1e-3. This model identifies the surgical tools’ presence within the frames extracted from our video data. We select the frames in which the count of detected tools matches the number of provided labels. This filtering process yielded a refined dataset comprising 29,562 frames for training and 904 frames for validation.
3) To increase the dataset, we use a form of pseudo-labeling. Initially, we train a classification model using the filtered dataset. This model uses the same encoder architecture described in the previous section, with the inclusion of a classification head. Additionally, we maintain the same loss function and optimizer configurations. After the training phase, we used the classification model to refine the labels for frames where the number of detected tools did not match the provided labels. We accomplish this by removing labels with the lowest probability, which contributes to the refinement of our final dataset. As a result, our final dataset includes all 47,227 training images and 1,073 validation images.
Detection label refinement
4) To transition from classification to detection we first create a segmentation network similar to the one described in step (2), however, at the end of the encoder we add a classification head. Initially, we train the encoder with a classification head on the entire classification dataset. Subsequently, we generate gradient-based saliency maps from the encoder and filter them using the output of the segmentation network described in step (2), solely using the prediction mask of the tool-clevis, or tools-tip for the specified tools in the data description.
5) While we acknowledge the potential presence of noise in our pseudo-segmentation masks, we have taken steps to mitigate its impact. Specifically, we adopt a symmetrical cross-entropy loss function proposed by Wang et al. [94], which enhances the robustness of our segmentation model against label noise. The total loss function during this phase is an equally weighted combination of the classification loss (cross-entropy) and segmentation loss (symmetrical cross-entropy). We retain the Adam optimizer with a learning rate of 1e-3 for consistency.
6) Additionally, we assume that, with several iterations, the output of our segmentation network would surpass the quality of the pseudo masks. Consequently, we apply a simple form of self-training, often employed in semi-supervised learning methods [95]. After 5, we computed a segmentation loss (symmetrical cross-entropy) for both the pseudo-segmentation masks generated in step (4) and the output of the segmentation network. The final loss function is as follows:
$\mathcal{L}_{\text{total}}=\alpha\mathcal{L}_{\text{cls}}+(\beta\mathcal{L}_{\text{seg}_{\text{masks}}}+(1-\beta)\mathcal{L}_{\text{seg}_{\text{output}}})$ (1)
Here the initial value of $\alpha$ and $\beta$ is 1 and decreases every epoch with 0.05. Causing the weight of the loss calculated based on the segmentation output to increase.
Final detection model
7) After obtaining predictions from the final segmentation model, a filtering process is applied to retain only the largest areas corresponding to the image classification labels, while eliminating smaller regions. Subsequently, the processed masks are transformed into a Yolo format, which serves as the input for training our YoloV8 detection model.
4.9.2 Model Training
The segmentation model, as described in (2), undergoes training for 25 epochs with a batch size of 4, utilizing images of resolution 512x512 pixels. Following this, the classification model outlined in (3) is initially trained for 10 epochs on the primary datasets, employing the same batch size and image resolution.
The encoder component of the segmentation model, detailed in (4), (5), and (6), is first trained on the complete classification dataset, using identical hyperparameters as the previous step. Subsequently, saliency masks are generated, and the entire encoder-decoder structure undergoes training for 5 epochs. For the subsequent 20 epochs, the model is trained utilizing the loss function described in Equation 1.
The final YOLOv8 detection model is trained for 25 epochs, with a batch size of 4. All models utilize ImageNet-initialized weights, and experiments are conducted on an RTX 3090Ti GPU (NVIDIA Corp., CA, USA).
4.9.3 Preliminary Performance
This method eventually resulted in a mean average precision of 0.0934 on the final test set. While our current method has shown promise in localizing surgical tools within individual frames, it does not yet harness the valuable temporal information present in video sequences. An immediate enhancement could involve integrating a tracking algorithm in conjunction with our approach, enabling more robust tool tracking over time.
Nevertheless, one of the standout advantages of our methodology lies in its scalability. This scalability paves the way for future developments with even larger datasets. Currently, we have limited the data input to just 2 frames per video. However, with additional resources and computational power, we have the potential to significantly increase this number without relying on additional human inference.
4.10 Results
Challenge evaluations were conducted on a private, concealed test set comprising individual surgical video files at 1 FPS. The Grand Challenge automated algorithm submission and evaluation system generated the results. Participants did not have access to the test data and labels. The evaluation of submissions employed the standard COCO dataset bounding box detection metric to determine the winners. The COCO (Common Objects in Context) dataset bounding box detection metric is a widely used evaluation measure in computer vision, particularly in the field of object detection. The bounding box detection metric evaluates the accuracy of object localization and classification by comparing the predicted bounding boxes of objects in each frame to the ground truth bounding boxes provided by the dataset. The metric is expressed here as mAP (mean Average Precision) at various Intersection over Union (IoU) thresholds. The IoU measures the overlap between the predicted bounding box and the ground truth bounding box, and IoU thresholds range from 0.5 to 0.95, with a step size of 0.05. Table 21 is a summary of the challenge results.
Rank Team mAP
1st PUMCH 0.4669
2nd ZJURealdoctor 0.3899
3rd Jmees 0.3478
4th seventeen 0.3404
5th ANL-Surg 0.3377
6th SDS-HD 0.3074
7th HVRL 0.1774
8th LozaGera 0.1688
9th CAIR-HK 0.1489
10th AIT 0.1139
11th TUE-VCA 0.0934
12th MapleLab 0.0007
13th LabRen-CUHK 0
Table 21: Highest Performance Results by Team
There is a significant variation in performance among the teams, as indicated by the range of mAP scores. A common characteristic among these top-performing teams is the generation of pseudo-labels to aid the training process. For example, PUMCH employed OSTrack for object tracking based on the provided annotated frames, ANL-Surg performed backward and forward tracking using Mixformer, and SDS-HD developed a semi-automatic annotation pipeline. All of these teams utilized attention-based models at some point.
4.11 Discussion
Considering the team’s results as a whole, some general observations can be made. As with the previous year, training the models with the noisy labels has proved to be challenging. Most successful strategies employed pseudo-labels to aid the training process. Many teams also made use of small manually annotated data sets, either labeled themselves, or by a previous year’s team. These labeled data sets where then bootstrapped to created larger labeled data sets. And, importantly, many teams followed approaches laid out by teams from the prior year.
Figure 45: Overview of challenge results from 2022 to 2023 demonstrating a robust increase in model performance. Results from the 2022 and 2023 challenge are plotted alongside the 2023 models evaluated on the 2022 test set.
To provide more context for this year’s results we compared them alongside those from the identical challenge of 2022 (Category 2). On the whole the models clearly performed better (compare the left and right bar plots, Fig 45). However, the 2023 test set was different from that of 2022. To control for this we evaluated this year’s models on the 2022 test set (middle bar plots, Fig 45). As can be seen from the overall decreases in performance, the 2022 test set was more challenging. However, despite this the 2023 models still outperformed those from the previous year. This is general trend of year-over-year improvements is precisely what the challenge was designed to foster, and why we intend to continue with the challenge in the future.
The current state-of-the-art models for object detection, such as CO-DETR [96], have achieved remarkable mAP values of up to 0.66 when trained and evaluated on general datasets such as COCO. Comparing these high-performance models with the those of the teams participating in this challenge reveals an apparent performance gap. This disparity can likely be attributed to many factors, but chief among them may be the specialized nature of surgical tool detection, which presents unique challenges not encountered in general object detection tasks. While our dataset contains thousands of video clips, the individual frames are often very similar and highly correlated, due to the nature of the surgical field (e.g. a background of red colored organs and metallic tools in the foreground). When contrasting this dataset with images from the COCO dataset, which may contain passenger jets or images of flowers, the differences are readily apparent. As such this dataset demands a relatively high degree of exactitude when distinguishing categorical labels. Closing the performance gap between specialized tasks like surgical tool detection and general object detection remains an ongoing challenge, with implications for improving medical imaging and healthcare.
These complications notwithstanding, the efforts of the participating teams demonstrate valuable progress in this domain and lay the groundwork for future advancements. The diverse range of teams participating in the challenge, drawn from institutions around the world, showcases the competitive landscape of surgical tool detection research, and a cause for an optimistic outlook.
5 Conclusion
Recent years have produced remarkable advancements in the area of computer vision. This is especially true with many well established public datasets. While the same performance cannot be claimed for similar problems in surgical data science, significant gains are being made. Future work may determine that detection in surgical images is quantifiably more difficult than in other typical image datasets (e.g. COCO, ImageNet). Regardless, we will continue to host this competition since we believe the ability to classify and segment images in these highly unstructured settings is crucial for surgical data science, and the field of computer vision in general.